### v1

# Efficient Neural Common Neighbor for Temporal Graph Link Prediction

Abstract
Temporal graphs are ubiquitous in real-world scenarios, such as social network, trade and transportation. Predicting dynamic links between nodes in a temporal graph is of vital importance. Traditional methods usually leverage the temporal neighborhood of interaction history to generate node embeddings first and then aggregate the source and target node embeddings to predict the link. However, such methods focus on learning individual node representations, but overlook the pairwise representation learning nature of link prediction and fail to capture the important pairwise features of links such as common neighbors (CN). Motivated by the success of Neural Common Neighbor (NCN) for static graph link prediction, we propose TNCN, a temporal version of NCN for link prediction in temporal graphs. TNCN dynamically updates a temporal neighbor dictionary for each node, and utilizes multi-hop common neighbors between the source and target node to learn a more effective pairwise representation. We validate our model on five large-scale real-world datasets from the Temporal Graph Benchmark (TGB), and find that it achieves new state-of-the-art performance on three of them. Additionally, TNCN demonstrates excellent scalability on large datasets, outperforming popular GNN baselines by up to 6.4 times in speed. Our code is available at https://github.com/GraphPKU/TNCN.
1 Introduction
Temporal graphs are increasingly utilized in contemporary real-world applications. Complex systems such as social networks (Yang et al., 2017; Min et al., 2021; Nguyen et al., 2017), trade and transaction networks (Zhang et al., 2018; Yan et al., 2021), and recommendation systems (Wu et al., 2022; Yin et al., 2019) are prime examples. These systems evolve dynamically over time, exhibiting different characteristics at various temporal moments. Recently, there has been a marked increase in the representation learning tasks on temporal graphs in the graph research community. At the same time, a prominent graph learning tool, Graph Neural Network (GNN) (Scarselli et al., 2008), has been developed to model node, link, and graph tasks. GNNs generally learn node embeddings by iteratively aggregating embeddings from neighboring nodes. They have demonstrated exceptional performance in graph representation learning, achieving state-of-the-art results across numerous tasks.
There remains a significant gap between static graphs and the increasingly prevalent temporal graphs. Temporal graphs incorporate discrete or continuous timestamps attached to edges, providing a more precise depiction of the graph evolution process. Meanwhile, many methodologies for static graphs are not applicable due to the additional constraints imposed by timestamps on causality. One can only utilize nodes and edges that precede a given time, implicitly resulting in numerous graph instances to analyze. Building on the success of time sequence modeling, Kumar et al. (2019); Trivedi et al. (2019); Rossi et al. (2020) propose memory-based temporal graph networks aimed at learning both short- and long-term dependencies. These methods, particularly Temporal Graph Networks (TGN), have achieved notable success in temporal node classification. Additionally, Transformer-based models, such as those proposed by Wang et al. (2021a); Xu et al. (2020), employ the multi-head attention mechanism to capture both cross and internal relationships within a temporal graph.
(a) a temporal graph
(b) computation tree of $u$ and $v$
(c) common neighbor method
Figure 1: Figure (a) shows a failure case of link prediction based on node-wise representation learning. Such methods cannot distinguish node $u$ and $v$ because they possess the same temporal computation tree in figure (b), thus generating the same node representation. However, when we try to learn their pair-wise representation, i.e. $(u,w)$ and $(v,w)$, we can observe that $v$ has a temporal common neighbor $b$ with node $w$ while $u$ doesn’t, as shown in figure (c). Thus with the same computation graph, we only need to utilize the extra node $b$’s embedding to distinguish $(u,w)$ and $(v,w)$.
However, these models may exhibit inherent flaws when addressing link prediction tasks in graphs. Such tasks require the model to predict the existence of a target link given specific prerequisite information. The aforementioned approaches primarily focus on node-wise representation learning, using the node embeddings of source and destination nodes to predict the existence of a target link. While straightforward, this method often fails to precisely capture the relationship between two nodes and other more complex graph structure features. For example, in Figure 1, the node-wise method might struggle to predict whether node $w$ prefers to interact with node $u$ or $v$. Generally speaking, these models are constrained by their focus on node-level representation, limiting their ability to capture broader contexts and higher-order patterns.
Considering these shortcomings, Zhang and Chen (2018) highlight the importance of pair-wise representation and propose labeling trick to mark the source and destination nodes when learning their pair-wise representation by GNN. Labeling trick (Zhang et al., 2021) is shown to greatly enhance GNN performance on static graph link prediction. Building on this, Wang et al. (2021c); Luo and Li (2022); Yu et al. (2023) develop models that leverage information from temporal surrounding nodes, successfully extending the pair-wise learning approach to dynamic graphs. They extract features using decreasing-timestamp random walks or joint neighborhood structures to generate multi-level embeddings for the center nodes, facilitating node classification and link prediction with pair-wise representations. However, these graph-based models often incur significant computational and memory costs due to extracting temporal neighborhoods and applying message passing on them for each node/link to predict, hindering their application in real-world, large-scale scenarios.
In general, graph-based representation learning methods can significantly enhance model capabilities, yet their high computational cost limits widespread application. For example, the Temporal Graph Benchmark (TGB) (Huang et al., 2023) contains many high-quality real-world temporal graphs. However, most graph-based models choose to only evaluate on a subset of small graphs due to their unaffordable cost of training and evaluation on large-scale datasets. On the contrary, memory-based models (Rossi et al., 2020) update node representations sequentially following the event stream, thus is significantly faster than graph-based methods yet might lose important graph information. Motivated by these observations, we propose an expressive but efficient model, Temporal Neural Common Neighbor (TNCN), to combine the merits of both. By integrating a memory-based backbone with Neural Common Neighbor (Wang et al., 2023), TNCN learns expressive pairwise representations while maintaining high efficiency akin to memory-based models. Consequently, TNCN is suitable for large-scale temporal graph link prediction.
We conducted experiments on five large-scale real-world temporal graph datasets from TGB. TNCN achieved new state-of-the-art results on three datasets and overall ranked first compared to 10 competitive baselines, demonstrating its effectiveness. To examine its scalability, we selected datasets with temporal edges ranging from $\mathcal{O}(10^{5})$ to $\mathcal{O}(10^{7})$ and node numbers ranging from thousands to millions. We found that TNCN achieved 1.9$\sim$4.7x speedup in training and 2.1$\sim$6.4x speedup in inference compared to graph-based models, while maintaining a similar scale of time consumption as memory-based models.
2 Related Work
2.1 Memory-based Temporal Graph Representation Learning
Temporal graph learning has garnered significant attention in recent years. A classic approach in this domain involves learning node memory using continuous events with non-decreasing timestamps. Kumar et al. (2019) propose a coupled recurrent neural network model named JODIE that learns the embedding trajectories of users and items. Another contemporary work DyRep (Trivedi et al., 2019) aims to efficiently produce low-dimensional node embeddings to capture the communication and association in dynamic graphs. Rossi et al. (2020) introduces a memory-based temporal neural network known as TGN, which incorporates a memory module to store temporal node representations updated with messages generated from the given event stream. Apan (Wang et al., 2021b) advances the methodology by integrating asynchronous propagation techniques, markedly increasing the efficiency of handling large-scale graph queries. EDGE (Chen et al., 2021) emerges as a computational framework focusing on increasing the parallelizability by dividing some intermediate nodes in long streams each into two independent nodes while adding back their dependency by training loss. Chen et al. (2023) extend the update method for the node memory module, introducing an additional hidden state to record previous changes in neighbors. Complementing these efforts, additional contributions such as Edgebank (Poursafaei et al., 2022) and DistTGL (Zhou et al., 2023) have been directed towards formalizing and accelerating memory-based temporal graph learning methods.
2.2 Graph-based Temporal Graph Representation Learning
Subsequent works have incorporated the temporal neighborhood structure into temporal graph learning. CAWN (Wang et al., 2021c) employs random anonymous walks to model the neighborhood structure. TCL (Wang et al., 2021a) samples a temporal dependency interaction graph that contains a sequence of temporally cascaded chronological interactions. TGAT (Xu et al., 2020) considers the temporal neighborhood and feeds the features into a temporal graph attention layer utilizing a masked self-attention mechanism. NAT (Luo and Li, 2022) constructs a multi-hop neighboring node dictionary to extract joint neighborhood features and uses a recurrent neural network (RNN) to recursively update the central node’s embedding. This information is then processed by a neural network-based encoder to predict the target link. DyGFormer (Yu et al., 2023), instead, leverages one-hop neighbor embeddings and the co-occurrence of neighbors to generate features, which are well-patched and subsequently fed into a Transformer (Vaswani et al., 2017) decoder to obtain the final prediction.
2.3 Link Prediction Methods
Link prediction is a fundamental task in graph analysis, aiming to determine the likelihood of a connection between two nodes. Early investigations posited that nodes with greater similarity tend to be connected, which led to a series of heuristic algorithms such as Common Neighbors, Katz Index, and PageRank (Newman, 2001; Katz, 1953; Page et al., 1999). With the advent of GNNs, numerous methods have attempted to utilize vanilla GNNs for enhancing link prediction, revealing sub-optimal performance due to the inability to capture important pair-wise patterns such as common neighbors (Zhang and Chen, 2018; Zhang et al., 2021; Liang et al., 2022). Subsequent research has focused on infusing various forms of inductive biases to retrieve intricate pair-wise relationships. For instance, SEAL (Zhang and Chen, 2018), Neo-GNN (Yun et al., 2021), and NCN (Wang et al., 2023) have integrated neighbor-overlapping information into their design. BUDDY (Chamberlain et al., 2022) and NBFNet (Zhu et al., 2021) have concentrated on extracting higher-order structural information. Additionally, Mao et al. (2023); Li et al. (2024) have contributed to a more unified framework encompassing different heuristics.
3 Preliminaries
Definition 3.1.
(Temporal Graph) Temporal graph can be typically categorized into two kinds, discrete-time (DTDG) and continuous-time (CTDG) dynamic graph. While DTDG can be represented as a special sequence of graph snapshots in the form of $\mathcal{G}=\{(\mathcal{G}_{1},t_{1}),(\mathcal{G}_{2},t_{2}),\cdots,(\mathcal{G}_{N},t_{N})\}$, it can always be transformed into its corresponding form in CTDG. So here we mainly focus on continuous time temporal graph. We usually represent a CTDG as a sequence of interaction events: $\mathcal{G}=\{(u_{1},v_{1},t_{1}),(u_{2},v_{2},t_{2}),\cdots,(u_{n},v_{n},t_{n})\}$, where $u,v$ stand for source and destination nodes and $\{t_{i}\}$ are chronologically non-decreasing. We use $\mathcal{V}$ to denote the entire node set, and $\mathcal{E}$ the entire edge set. Note that each node or edge can be attributed, that is, there may be node feature $x_{u}$ for $u$ or edge feature $e_{u,v}^{t}$ attached to the event $(u,v,t)$.
Definition 3.2.
(Problem Formulation) Given the events before time $t^{*}$, i.e. $\{(u,v,t)\ |\ \forall\ t<t^{*}\}$, a link prediction task is to predict whether two specified node $u^{*}$ and $v^{*}$ are connected at time $t^{*}$.
Definition 3.3.
(Temporal Neighborhood) Given the center node $u$, the $k$-hop temporal neighbor set $(k\geq 0)$ before time $t$ is defined as $N_{k}^{t}(u)=\{v\ |\ \text{SPD}^{t}(u,v)=k\}$, where $\text{SPD}^{t}(u,v)$ refers to the shortest path distance using edges before time $t$. Then the $k$-hop common neighbors between node $u$ and $v$ are defined as $\text{CN}_{k}^{t}(u,v)=\mathop{\cup}\limits_{2(k-1)<k_{1}+k_{2}\leq 2k,k_{1},k_{2}\leq k}(N_{k_{1}}^{t}(u)\cap N_{k_{2}}^{t}(v))$, i.e., all nodes on at least one $2k$-distance path between $u,v$. Note that for $k=0$, we define the 00-hop common neighbor of $u$ and $v$ as $\text{CN}_{0}^{t}(u,v)=\{u,v\}$ if there exists at least one edge between $u$ and $v$ before time $t$, and $\emptyset$ otherwise. Finally, the $K$-hop temporal neighborhood is defined as $\mathop{\cup}\limits_{k=0}^{K}N_{k}^{t}(u)$.
4 Methodology
In this section, we introduce our Temporal Neural Common Neighbor (TNCN) model. TNCN comprises several key modules: the Memory Module, the Temporal CN Extractor, and the NCN-based Prediction Head. Special attention is given to the Temporal CN Extractor, designed to efficiently extract temporal neighboring structures and obtain multi-hop common neighbor information. The pipeline of TNCN is illustrated in Figure 2.
Figure 2: Pipeline of TNCN. TNCN operates through a sequential update and prediction framework that processes successive batches of messages. During the update phase, TNCN updates the neighbor dictionary and the node memory representations. In the prediction phase, the model retrieves neighbors to identify common neighbors, thereafter leveraging the representations of the target nodes and their common neighbors for prediction.
4.1 Memory Module
The memory module stores node memory representations up to time $t$. When a new event occurs, the memory of the source and destination nodes is updated with the message produced by the event. The computation can generally be represented by the following formulas:
$msg_{src}^{t}(u,v)=\textit{msgfunc}_{src}(e^{t}_{u,v}),$ $msg_{dst}^{t}(u,v)=\textit{msgfunc}_{dst}(e^{t}_{u,v}).$ (1)
$mem^{t}_{u}=\textit{upd}_{src}(mem^{t-}_{u},msg_{src}^{t}(u,v)),$ $mem^{t}_{v}=\textit{upd}_{dst}(mem^{t-}_{v},msg_{dst}^{t}(u,v)),$
where $mem^{t-}_{u}$ stands for the embedding of node $u$ before time $t$. Here, msgfunc is a learnable function, such as a linear projection or a simple identity function. Note that for different edge directions, i.e., from source to destination and vice versa, the msgfunc and updfunc can be learned separately. The memory module aids the model in managing both long-term and short-term dependencies, thereby reducing the likelihood of forgetting. During training and inference, node memory evolves dynamically as events occur. Updates to this module reflect the dynamic nature of the temporal graph.
4.2 Temporal CN Extractor
Our Temporal CN Extractor can efficiently perform multi-hop common neighbor extraction and aggregate their neural embeddings.
Extended Common Neighbor.
The definition of multi-hop common neighbors (CN) is given in Definition 3.3, extending the traditional 1-hop CN (i.e., nodes on 2-paths between $u$ and $v$) to $k$-hop CN (nodes on $2k$-paths between $u$ and $v$). Additionally, we define zero-hop CN to be $u$ and $v$ themselves, i.e., if there is at least one previous event between node $u$ and node $v$, then $u$ and $v$ are considered each other’s 0-hop common neighbors. The 0-hop CN not only records the historical interactions between two nodes, but also reveals the frequency of their interactions.
Efficient CN Extractor.
The CN Extractor is a crucial component of the TNCN model, contributing significantly to its high performance. It can efficiently gather pertinent information about a given center node and extract multi-hop common neighbors for a source-destination pair.
For each relevant node $u$, the extractor stores its historical interactions with other nodes as both source and destination. After a batch of events is processed by the model, the storage is updated with the latest interactions. This allows us to maintain a record of all historical interactions up to a certain timestamp, effectively constructing a dynamic lookup dictionary for fast retrieval during subsequent inference. To conserve memory, we can opt to save only the most recent events and relevant nodes.
During the inference stage, TNCN extracts temporal neighborhood information, i.e., historical interactions related to the query node batch. This information is organized as a Sparse Tensor, on which we perform self-multiplication to generate high-order adjacency connectivity. We use sparse tensor operations to obtain multi-hop common neighbors (CNs) and their embeddings, separating CNs from different hops. To extract the 0-hop common neighbors mentioned above in a unified way, we introduce an additional self-loop for each related node. This operation allows us to use the same tensor multiplications to obtain 0-hop CNs.
The utilization of Multi-hop Common Neighbors significantly boosts TNCN’s performance, resulting in higher scores in temporal link prediction tasks. Furthermore, by employing sparse tensors, our model achieves substantial reductions in both storage requirements and computational complexity, thereby decreasing time consumption and enhancing efficiency.
4.3 NCN-based Prediction Head
For source and destination nodes, we perform an element-wise product:
$X^{t}_{u,v}=emb_{u}^{t}\otimes emb_{v}^{t}.$ (2)
For multi-hop CN nodes, we aggregate their embeddings in each hop with sum pooling:
$NCN_{k}(u,v)=\mathop{\oplus}\limits_{w\in\text{CN}^{t}_{k}(u,v)}emb_{w}^{t}.$ (3)
These embeddings are then concatenated as the final pair-wise representation:
$repr(u,v)$ $=[X^{t}_{u,v}\ ||\ (\mathop{||}\limits_{k=0}^{K})NCN_{k}(u,v)].$ (4)
In the above, $\otimes$, $\oplus$, and ||||| | stand for element-wise product, element-wise summation, and concatenation of vectors, respectively. The pair-wise representation $repr(u,v)$ for nodes $u$ and $v$ is fed to a projection head to output the final link prediction.
5 Efficiency and Effectiveness of TNCN
In this section, we explore the two principal benefits of TNCN: efficiency and effectiveness. These advantages are demonstrated through an analysis of two core components within the framework for temporal graph link prediction: graph representation learning and link prediction methods. We categorize graph representation learning modules into two types: memory-based and $k$-hop-subgraph-based, according to the temporal scope of evolved events. Memory-based modules exhibit superior time complexity while maintaining good expressiveness in some situations, striking a balance between efficiency and performance. Furthermore, we highlight the deficiencies of existing link prediction methods on temporal graph learning and introduce the extended common neighbor approach. This method serves as a complementary addition for learning pair-wise representations while eliminating the necessity for message passing on entire graphs. Both our graph representation learning and link prediction techniques are designed with a unified optimization objective: to avoid message passing on entire subgraphs in favor of non-repetitive operations, culminating in a cohesive solution that is both efficient and effective.
5.1 Graph Representation Learning
Graph representation learning aims to develop an embedding function, denoted as $Emb$, which learns an embedding for each node encoding its structural and feature information within the graph. Specifically, given a new event represented as $(u,v,t)$, the function $Emb$ leverages prior events to generate meaningful embeddings. Approaches in this domain diverge in their handling of temporal dynamics; some opt to maintain a dynamic embedding for each node that is incrementally updated with each new event, while others choose to recalculate node embeddings by considering the entire historical context of events, thereby providing a more comprehensive reflection of past interactions. We classify these methodologies into two distinct types based on their operational mechanisms.
Definition 5.1.
Memory-based approach. Given a new event $(u,v,t)$, if $Emb$ conforms to the following form, the method is referred to as a memory-based approach.
$Emb(u,t)=f_{\textit{emb}}(Mem(u,t^{\prime})),~{}$ $Mem(u,t)=f_{\textit{mem}}(Mem(u,t^{\prime}),Mem(v,t^{\prime}),e_{u,v}^{t},t-t^{\prime}),$ (5)
$Emb(v,t)=f_{\textit{emb}}(Mem(v,t^{\prime})),~{}$ $Mem(v,t)=f_{\textit{mem}}(Mem(v,t^{\prime}),Mem(u,t^{\prime}),e_{v,u}^{t},t-t^{\prime}),$
where $f_{\textit{emb}}$ and $f_{\textit{mem}}$ are two learnable functions.
Definition 5.2.
$k$-hop-subgraph-based approach. Given a new event $(u,v,t)$, if $Emb$ conforms to the following form, the method is defined as a subgraph-based approach.
$Emb(u,t)=f_{\textit{emb}}(\mathcal{G}_{u,<t}^{k}),~{}~{}~{}~{}Emb(v,t)=f_{\textit{emb}}(\mathcal{G}_{v,<t}^{k}),$ (6)
where $\mathcal{G}_{u,<t}^{k}$ is a subgraph induced from $\mathcal{G}$ by node $u$’s $k$-hop temporal neighborhood $\mathop{\cup}\limits_{k=0}^{K}N_{k}^{t}(u)$, containing only the edges (events) with time $t^{\prime}<t$, and $f_{\textit{emb}}$ is a learnable function.
Effectiveness.
The analysis begins by assessing the effectiveness of the two paradigms. To do so, we first introduce the concept of $k$-hop event (LOV´ASZ et al., 1993).
Definition 5.3.
($k$-hop event & monotone $k$-hop event) A $k$-hop event is a sequence of consecutive edges $\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$ connecting the initial node $u_{0}$ to the final node $u_{k}$. For example, $\{(u,x,t^{\prime}),(x,v,t)\}$ is a $2$-hop event. In the case where $k=1$, the $k$-hop event reduces to a single interaction $(u,v,t)$. A monotone $k$-hop event is a $k$-hop event in which the sequence of timestamps $\{t_{u_{i},u_{i+1}}\mid i\in\{0,\ldots,k-1\},k\geq 1\}$ is strictly monotonically increasing.
Then, we analyze the expressiveness of the two approaches in terms of encoding $k$-hop event.
Theorem 5.4.
(Ability to encode $k$-hop events). Given a $k$-hop event $\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$, if the node embedding of $u_{0}$ at time $t_{u_{0},u_{1}}$ can be reversely recovered from the encoding $\textit{Enc}(\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\})$, then we say the encoding function Enc is capable of encoding the $k$-hop event. The following results outline the encoding capabilities of different learning paradigms:
  • Memory-based approaches can encode any $k$-hop events with $k=1$.
  • Memory-based approaches can encode any monotone $k$-hop events with arbitrary $k$.
  • $k$-hop-subgraph-based approaches can encode any $k^{\prime}$-hop events with $k^{\prime}\leq k$.
From Theorem 5.4, we can conclude that 1) memory-based approaches have superior expressiveness in encoding $k$-hop events compared to 1-hop-subgraph-based approaches, 2) memory-based approaches have superior expressiveness in encoding monotone $k$-hop events than $k^{\prime}$-hop-subgraph-based approaches when $k^{\prime}<k$, and 3) $k$-hop-subgraph-based approaches are not less expressive than memory-based approaches when $k$ is large enough.
While the memory-based approach does not consistently rival the expressiveness of the $k$-hop-subgraph paradigm, it possesses advantages in monotone events and long-history scenarios (where $k$-hop subgraphs would be unaffordable to extract).
Efficiency.
We then turn our attention to the efficiency of the two approaches. A pivotal factor is the frequency with which individual events are incorporated into computations. In memory-based approaches, each event is utilized a single time for learning, immediately following its associated prediction. Conversely, in the $k$-hop-subgraph-based method, an event may be employed multiple times, as it is revisited in different nodes’ temporal neighborhood and repeated been processed within each subgraph’s encoding (such as message passing) process. This discrepancy leads to divergent cumulative frequencies of event utilization throughout the learning process and results in the huge efficiency advantage of memory-based methods. We formalize this observation in the following:
Theorem 5.5.
(Learning method time complexity). Denote the time complexity of a learning method as a function of the total number of events processed during training. For a given graph $\mathcal{G}$ with the number of nodes designated as $|\mathcal{N}|$ and the number of edges as $|\mathcal{E}|$, the following assertions hold:
  • For memory-based approaches, the time complexity is $\Theta\left(|\mathcal{E}|\right)$.
  • For $k$-hop-subgraph-based approaches with $k=1$, the lower-bound time complexity is $\Omega\left(\frac{|\mathcal{E}|^{2}}{|\mathcal{N}|}\right)$, and the upper-bound time complexity is $\mathcal{O}\left(\frac{|\mathcal{E}|^{2}}{|\mathcal{N}|}+|\mathcal{E}||\mathcal{N}|\right)$.
  • For $k$-hop-subgraph-based approaches with $k=2$, the upper-bound time complexity is $\mathcal{O}\left(\left(\frac{|\mathcal{E}|^{2}}{|\mathcal{N}|}+|\mathcal{E}||\mathcal{N}|\right)^{\frac{3}{2}}\right)$.
Following Theorem 5.5, it becomes evident that the computational overhead incurred by a memory-based method is significantly lower than that of a $k$-hop-subgraph-based method, particularly as the value of $k$ increases. These results highlight the advantages of memory-based methods in mitigating the computational efficiency challenges associated with large-scale temporal graphs.
5.2 Link Prediction Technique
With the node embeddings obtained from the graph representation learning step, link prediction techniques involve aggregating the node embeddings in some way into link representations for link prediction. Most previous methods simply concatenate the source and destination node embeddings as the link representation $Emb(u,t)\mathop{||}Emb(v,t)$, losing a great amount of pairwise features (as illustrated in Figure 1). In order to use labeling trick, an alternative approach involves extracting a separate subgraph for each link to predict, which leads to good results but also significantly increases the complexity. TNCN, on the contrary, utilizes neural common neighbors as a decoupled and flexible solution. This approach can be seamlessly integrated into existing memory-based methods with minimal computational overhead, while retaining important structural information for link prediction.
Effectiveness.
In the following, we first demonstrate the effectiveness of TNCN by showing that it can capture three important pairwise features commonly used as effective link prediction heuristics, namely Common Neighbors (CN), Resource Allocation (RA), and Adamic-Adar (AA) (Newman, 2001; Adamic and Adar, 2003; Zhou et al., 2009), borrowing from NCN (Wang et al., 2023).
Theorem 5.6.
TNCN is strictly more expressive than CN, RA, and AA.
Essentially, TNCN uses node memory to substitute the constant or degree-based values in the three features, and the memory update scheme is sufficient to learn such values. In comparison, an approach that merely concatenates node-wise representations proves inadequate in capturing these heuristics.
Efficiency.
We then address the efficiency of TNCN. The computation of TNCN can be divided into two primary components: the generation of neighbor nodes’ embeddings and the execution of common neighbor lookups. Concerning the former, the memory-based approach intrinsically tracks all node embeddings, thus obviating the need for re-computation. As for the latter, we implement a fast common neighbor search algorithm by leveraging a sparse matrix structure that supports batch operations. Collectively, these factors contribute to the minimal additional overhead of TNCN compared to memory-based methods.
6 Experiments
This section assesses TNCN’s effectiveness and efficiency by answering the following questions:
Q1: What is the performance of TNCN compared with state-of-the-art baselines?
Q2: What is the computational efficiency of TNCN in terms of time consumption?
Q3: Do the extended common neighbors bring benefits to original common neighbors?
6.1 Experimental Settings
Datasets
We evaluate our model on five large-scale real-world datasets for temporal link prediction from the Temporal Graph Benchmark (Huang et al., 2023). These datasets span several distinct fields: co-editing network on Wikipedia, Amazon product review network, cryptocurrency transactions, directed reply network of Reddit, and crowdsourced international flight network. They vary in scales and time spans. Additional details about the datasets are provided in Appendix A. We set the evaluation metric as Mean Reciprocal Rank (MRR) consistent with the TGB official leaderboard.
Baselines
We systematically evaluate our proposed model against a diverse set of baselines known for their strong capacity to represent temporal graph dynamics. These include: a heuristic algorithm Edgebank (Yu et al., 2023), memory-based models DyRep (Trivedi et al., 2019) and TGN (Rossi et al., 2020) that obviate the need for frequent temporal subgraph sampling, and GraphMixer (Cong et al., 2023) which primarily employs an MLP-mixer. Additionally, we include various graph-based models such as CAWN (Wang et al., 2021c), NAT (Luo and Li, 2022), DyGFormer (Yu et al., 2023), TGAT (Xu et al., 2020), and TCL (Wang et al., 2021a), which learn from neighborhood structure information. To reconcile the trade-off between model efficacy and computational overhead, we evaluate two TNCN variants: the “TNCN-standard” (TNCN(S)), which incorporates up to 1-hop common neighbor information, and the more complex “TNCN-large” (TNCN(L)), which uses up to 2-hop common neighbors.
6.2 Experimental Results
Table 1: Test Performance of different models under MRR metric. The top three are emphasized by red, blue and black fonts. ‘NA’ denotes scenarios where a specific method was either not applied to the dataset or was unable to complete the validation and testing phases within a reasonable timeframe.
Model Wiki Review Coin Comment Flight
TGN 0.528 0.387 0.737 0.622 0.705
DyGformer 0.798 0.224 0.752 0.670 NA
NAT 0.749 0.341 NA NA NA
CAWN 0.711 0.193 NA NA NA
Graphmixer 0.118 0.521 NA NA NA
TGAT 0.141 0.355 NA NA NA
TCL 0.207 0.193 NA NA NA
DyRep 0.050 0.220 0.452 0.289 0.556
Edgebank(tw) 0.571 0.025 0.580 0.149 0.387
Edgebank(un) 0.495 0.023 0.359 0.129 0.167
TNCN-standard 0.720 0.298 0.739 0.692 0.803
TNCN-large 0.717 0.317 0.764 0.662 NA
Reply to Q1: TNCN possesses remarkable performance.
We conducted comprehensive evaluations of prevailing methods on the TGB. The main results are summarized in Table 1. It is evident from the table that TNCN attains new SOTA performance on three out of five datasets. Additionally, TNCN demonstrates competitive results on the remaining two datasets. TNCN almost consistently surpasses memory-based models such as TGN and DyRep, which can be attributed to its utilization of supplementary structural information. Even in comparison to powerful graph-based models, including NAT and DyGFormer, TNCN still matches or exceeds their performance, underscoring the effectiveness of its integration of node memory and pair-wise features. The only dataset where TNCN exhibits relatively poorer performance is the Review dataset. This can be ascribed to the dataset’s predominant reliance on feature embedding and its high "surprise" metric, wherein prior events have a diminished correlation with subsequent events, potentially reducing the impact of CNs.
Reply to Q2: TNCN shows great scalability on large datasets.
(a) tgbl-wiki (log scale)
(b) tgbl-review (linear scale)
Figure 3: Time Consumption of Memory and Graph-based Method on Wiki and Review Datasets
To evaluate computational efficiency, we collected the time consumption when applying all baselines to the Wiki and Review datasets, as depicted in Figure 3. Compared with memory-based methods, TNCN exhibits a comparable order of magnitude in terms of time consumption. However, when benchmarked against graph-based models, TNCN demonstrates a substantial acceleration, achieving approximately 1.9 to 4.7 times speedup during the training phase and a 2.1 to 6.4 times increase in inference speed. Notably, the scalability concerns become even more evident as the size of the dataset expands; several graph-based models cannot complete the validation and testing processes within a reasonable time budget. The primary factors contributing to TNCN’s efficiency are the synergistic, time-efficient design of its two core components and the implementation of the Efficient CN Extractor that facilitates batch operations through parallel processing.
Reply to Q3: Extended CN brings improvements.
Table 2: Test performance of TGN and TNCN with different ranges of common neighbors
Model Wiki Review Coin Comment
TGN 0.528 0.387 0.737 0.622
TNCN-1-hop-CN 0.621 0.374 0.737 0.641
TNCN-0$\sim$1-hop-CN 0.720 0.298 0.739 0.692
TNCN-0$\sim$2-hop-CN 0.717 0.317 0.764 0.662
To elucidate the benefits of extended CN, we conducted an ablation study on the hop range of common neighbors. The results are shown in Table 2. The conventional NCN methodology considers only 1-hop CN. However, this approach may not be universally applicable across all temporal network topologies. For instance, bipartite graphs lack such 1-hop CN in their structure, necessitating the consideration of 2-hop CN. Additionally, memory-based methodologies may omit a notable aspect: they generally find it difficult to quantify the frequency of interactions between a given pair of nodes, which brings the need for 0-hop CN.
[…]
7 Conclusion and Limitation
We propose TNCN for temporal graph link prediction, which employs a temporal common neighbor extractor combined with a memory-based node representation learning module. TNCN has achieved new state-of-the-art results on several real-world datasets while maintaining excellent scalability to handle large-scale temporal graphs.
However, based on our observation of TNCN’s performance on the TGBL-Review dataset, there are some limitations in our model. Specifically, datasets with high surprise values, such as the Review dataset, tend to make it more challenging for TNCN to accurately predict the probability of future connections. This indicates that while TNCN performs well overall, it may struggle with datasets that exhibit high variability or unexpected patterns. Further research is needed to address these challenges and improve the model’s robustness in such scenarios.
Appendix A Datasets
Table 3 shows some detailed datasets statistics of TGB and 4 shows several temporal graph datasets commonly used by previous work. Through the two tables we can observe that TGB official datasets possess temporal graphs with larger scale to 10 million, 10 times surpassing the largest previous datasets such as LastFM, Flight, Contact. With the aim to examine our TNCN model’s efficiency, we choose the increasingly accepted datasets TGB.
Table 3: TGB Dataset Statistics
Dataset Domain Nodes Edges Steps Surprise Edge Properties
tgbl-wiki interact 9,227 157,474 152,757 0.108 W: $\times$, Di: ✓, A: ✓
tgbl-review rating 352,637 4,873,540 6,865 0.987 W: ✓, Di: ✓, A: $\times$
tgbl-coin transact 638,486 22,809,486 1,295,720 0.120 W: ✓, Di: ✓, A: $\times$
tgbl-comment social 994,790 44,314,507 30,998,030 0.823 W: ✓, Di: ✓, A: ✓
tgbl-flight traffic 18143 67,169,570 1,385 0.024 W: $\times$, Di: ✓, A: ✓
Table 4: Previous Dataset Statistics
Datasets Domains Nodes Links NL Feat Bipartite Duration Unique Steps Time Granularity
Wikipedia Social 9,227 157,474 – & 172 $\checkmark$ 1 month 152,757 Unix timestamps
Reddit Social 10,984 672,447 – & 172 $\checkmark$ 1 month 669,065 Unix timestamps
MOOC Interaction 7,144 411,749 – & 4 $\checkmark$ 17 months 345,600 Unix timestamps
LastFM Interaction 1,980 1,293,103 – & – $\checkmark$ 1 month 1,283,614 Unix timestamps
Enron Social 184 125,235 – & – $\times$ 3 years 22,632 Unix timestamps
UCI Social 1,899 59,835 – & – $\times$ 196 days 58,911 Unix timestamps
Flights Transport 13,169 1,927,145 – & 1 $\times$ 4 months 122 days
UN Trade Economics 255 507,497 – & 1 $\times$ 32 years 32 years
Contact Proximity 692 2,426,279 – & 1 $\times$ 1 month 8,064 5 minutes
Appendix B TNCN Model Configuration
Network Choice
In our experiment, the changeable neural networks are chosen as follows:
In Memory Module, we choose $Identity$ as msgfunc and GRU as $upd$. In inference stage we process node memory with Graph Attention Embedding to get the temporal representation. As for Prediction Head, we finally choose $Linear$ as the $repr$ function.
Hyper-parameter
Several detailed hyper-parameters for TNCN are shown in table 5, which can help researchers to reproduce the experiment performance as reported in this paper. Additionally, we set $mem\_dim$, $emb\_dim$, $time\_dim$ 100, 100, 100 respectively.
Table 5: Some Experiment Hyper-parameters
Dataset num_neighbors num_epoch patience
Wiki 10 20 5
Review 10 5 2
Coin 10 3 2
Comment 10 3 1
Appendix C Proofs
In this section, we give proofs on theorems 5.4, 5.5 and 5.6.
We commence with a restatement of Theorem 5.4 for clarity:
Theorem C.1.
(Ability of encoding $k$-hop event). Given a $k$-hop event $\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$. If the node embedding of $u_{0}$ at time $t_{u_{0},u_{1}}$, can be formally derived by the encoding function $\textit{Enc}(\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\})$, then the learning method is considered capable of encoding the $k$-hop event. The following outline the encoding capabilities of different learning schemes:
  • Memory-based approach can encode any $k$-hop events with $k=1$.
  • Memory-based approach can encode any monotone $k$-hop events with arbitrary $k$.
  • $k$-hop-subgraph-based approach can encode any $k^{\prime}$-hop events with $k^{\prime}\leq k$
Proof.
[…]
We consider the encoding of an arbitrary monotonically increasing $k$-hop temporal event sequence within a memory-based approach. The induction principle is applied to demonstrate the capability of this approach. For the base case, $k=1$, the encoding has been shown to be feasible. Now, assume the proposition holds for a $k\textquoteright$ ’-hop event; that is, any $k\textquoteright$ ’-hop temporal sequence of monotonically increasing events can be encoded using a memory-based approach. This assumption implies that there exists an embedding function such that
[…]
$Emb(u_{0},t_{u_{0},u_{1}})$ $=f_{\textit{emb}}(Mem(u_{0},t_{u_{1},u_{2}}))$ (8)
$=f_{\textit{emb}}(f_{\textit{mem}}(Mem(u_{0},t_{u_{1},u_{2}}),Mem(u_{1},t_{u_{1},u_{2}}),e_{u_{0},u_{1}}^{t_{u_{0},u_{1}}},t_{u_{0},u_{1}}-t_{u_{1},u_{2}}),$
$Mem(u_{1},t_{u_{1},u_{2}})$ $=f_{\textit{emb}}^{-1}(Emb(u_{1},t_{u_{1},u_{2}}))$
$=f_{\textit{emb}}^{-1}(\textit{Enc}(\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{1,\ldots,k^{\prime}\},k^{\prime}\geq 1\})$
[…]
We consider a $k$-hop-subgraph-based approach for our analysis. It is evident that a $k$-hop subgraph encompasses any $k\textquoteright$ ’-hop events, where $k\textquoteright\leq k$. Furthermore, the aggregation methodology assimilates all nodes contained within the subgraph. Collectively, these observations substantiate the theorem in question.
We commence with a restatement of Theorem 5.5 for clarity:
Theorem C.2.
(Learning method time complexity). Denote the time complexity of a learning method as a function of the total number of events processed during training. For a given graph ${\mathcal{G}}$ with the number of nodes designated as $|{\mathcal{N}}|$ and the number of edges as $|{\mathcal{E}}|$, the following assertions hold:
  • For the memory-based approach, the time complexity is $\Theta\left(|{\mathcal{E}}|\right)$.
  • For $k$-hop-subgraph-based with $k=1$, the lower-bound time complexity is $\Omega\left(\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}\right)$, and the upper-bound time complexity is ${\mathcal{O}}\left(\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}+|{\mathcal{E}}||{\mathcal{N}}|\right)$
  • For $k$-hop-subgraph-based with $k=2$, the upper-bound time complexity is ${\mathcal{O}}\left((\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}+|{\mathcal{E}}||{\mathcal{N}}|)^{\frac{3}{2}}\right)$
Proof.
In the proposed theorem, the time complexity is denoted as the aggregate quantity of events processed throughout the training phase. The objective herein is to ascertain the precise count of such utilized events.
In the context of the memory-based methodology, it is evident that each event is utilized a singular time. Consequently, the cumulative number of events is expressed as $|\mathcal{E}|$, which infers that the time complexity adheres to the order of $\Theta\left(|\mathcal{E}|\right)$.
In the context of $k$-hop-subgraph-based algorithms wherein $k=1$, an event $(u,v,t)$ is exploited once for every incident event within the neighborhood of vertices $u$ or $v$. Without loss of generality, we focus on all events within the $1$-hop-subgraph of vertex $u$. The aggregate count of events processed is given by $\sum_{i=1}^{d(u)}i=\Theta\left(d(u)^{2}\right)$, where $d(u)$ denotes the degree of vertex $u$. Consequently, the computational complexity is fundamentally proportional to $\sum_{u\in{\mathcal{N}}}d(u)^{2}$. Drawing on the results of de Caen [1998], the lower bound on the time complexity is established as $\Omega\left(\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}\right)$, whereas the upper bound is determined as ${\mathcal{O}}\left(\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}+|{\mathcal{E}}||{\mathcal{N}}|\right)$
In the context of $k$-hop-subgraph-based algorithms wherein $k=2$,, we adopt similar strategy where each event$(u,v,t)$ will only be utilized once another event within the subgraph of $u$ or $v$ is firstly considered. The total number of events can be formulated as $\sum_{u\in{\mathcal{N}}}d(u)\sum_{v\in{\mathcal{N}}_{u}}\sum_{w\in{\mathcal{N}}_{v}}d(w)$. Replacing $d(u)$ as $X_{i}$, $\sum_{v\in{\mathcal{N}}_{u}}\sum_{w\in{\mathcal{N}}_{v}}$ as $Y_{i}$, we reformulated is as $\sum_{i\in|{\mathcal{N}}|}X_{i}Y_{i}$, satisfying $\sum_{i\in|{\mathcal{N}}|}X_{i}^{2}=\sum_{u\in{\mathcal{N}}}d(u)^{2}$ and $\sum_{i\in|{\mathcal{N}}|}Y_{i}^{2}=\sum_{u\in{\mathcal{N}}}d(u)^{4}$. Following Cauchy inequality and conclusions of $\sum_{u\in{\mathcal{N}}}d(u)^{2}$, we got the the upper-bound time complexity is ${\mathcal{O}}\left((\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}+|{\mathcal{E}}||{\mathcal{N}}|)^{\frac{3}{2}}\right)$
We commence with a restatement of Theorem 5.6 for clarity:
Theorem C.3.
TNCN is strictly more expressive than CN, RA, and AA.
We first give definitions of these structural features under temporal settings. Given two nodes $u$ and $v$, the structural features before time $t$ are defined as follows:
$CN(u,v,t)$ $=\sum_{w\in N_{1}^{t}(u)\cap N_{1}^{t}(v)}1,$ (9)
$RA(u,v,t)$ $=\sum_{w\in N_{1}^{t}(u)\cap N_{1}^{t}(v)}\frac{1}{d(w)},$
$AA(u,v,t)$ $=\sum_{w\in N_{1}^{t}(u)\cap N_{1}^{t}(v)}\frac{1}{\log d(w)}$
Given a node $u$, the degree of node $u$ is the number of events $e$ with an endpoint at node $u$. Without loss of generality (WLOG), we consider node $u$ as the source node, and the events are $\{(u,v_{i},t_{i})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$. Each time a new event is given, the embedding of node $u$ is updated by
$mem_{u}^{t_{i}}=\textit{upd}_{src}(mem_{u}^{t_{i-1}},\textit{msgfunc}_{src}(e_{u,v_{i}}^{t_{i}})).$ (10)
With the MPNN universal approximation theorem, msgfunc can be a constant function, and $\textit{upd}_{src}$ can be an addition function. Thus,
$mem_{u}^{t_{k-1}}=d(u).$ (11)
Then the embedding can learn arbitrary functions of node degrees, i.e.,
$emb_{u}^{t}=f(d(u)).$ (12)
Thus, the neural common neighbor $TNCN_{1}(u,v)=\mathop{\oplus}_{w\in N_{1}^{t}(u)\cap N_{1}^{t}(v)}emb_{w}^{t}$ can express Equation 9.
Extending to situations where the common neighbor node has some features we want to learn, the traditional CN, RA, and AA cannot accommodate this. However, our TNCN can express these features, demonstrating that TNCN is strictly more expressive than CN, RA, and AA.

### v2

@@
 Abstract
-Temporal graphs are ubiquitous in real-world scenarios, such as social network, trade and transportation.
-Predicting dynamic links between nodes in a temporal graph is of vital importance.
-Traditional methods usually leverage the temporal neighborhood of interaction history to generate node embeddings first and then aggregate the source and target node embeddings to predict the link.
-However, such methods focus on learning individual node representations, but overlook the pairwise representation learning nature of link prediction and fail to capture the important pairwise features of links such as common neighbors (CN).
-Motivated by the success of Neural Common Neighbor (NCN) for static graph link prediction, we propose TNCN, a temporal version of NCN for link prediction in temporal graphs.
-TNCN dynamically updates a temporal neighbor dictionary for each node, and utilizes multi-hop common neighbors between the source and target node to learn a more effective pairwise representation.
-We validate our model on five large-scale real-world datasets from the Temporal Graph Benchmark (TGB), and find that it achieves new state-of-the-art performance on three of them.
-Additionally, TNCN demonstrates excellent scalability on large datasets, outperforming popular GNN baselines by up to 6.4 times in speed.
+Temporal graphs are widespread in real-world applications such as social networks, as well as trade and transportation networks.
+Predicting dynamic links within these evolving graphs is a key problem.
+Many memory-based methods use temporal interaction histories to generate node embeddings, which are then combined to predict links.
+However, these approaches primarily focus on individual node representations, often overlooking the inherently pairwise nature of link prediction.
+While some recent methods attempt to capture pairwise features, they tend to be limited by high computational complexity arising from repeated embedding calculations, making them unsuitable for large-scale datasets like the Temporal Graph Benchmark (TGB).
+To address the critical need for models that combine strong expressive power with high computational efficiency for link prediction on large temporal graphs, we propose Temporal Neural Common Neighbor (TNCN).
+Our model achieves this balance by adapting the powerful pairwise modeling principles of Neural Common Neighbor (NCN) to an efficient temporal architecture.
+TNCN improves upon NCN by efficiently preserving and updating temporal neighbor dictionaries for each node and by using multi-hop common neighbors to learn more expressive pairwise representations.
+TNCN achieves new state-of-the-art performance on Review from five large-scale real-world TGB datasets, 6 out of 7 datasets in the transductive setting and 3 out of 7 in the inductive setting on small- to medium-scale datasets.
+Additionally, TNCN demonstrates excellent scalability, outperforming prominent GNN baselines by up to 30.3 times in speed on large datasets.
 Our code is available at https: //github.com/GraphPKU/TNCN.
 1 Introduction
-Temporal graphs are increasingly utilized in contemporary real-world applications.
-Complex systems such as social networks (Yang et al., 2017; Min et al., 2021; Nguyen et al., 2017), trade and transaction networks (Zhang et al., 2018; Yan et al., 2021), and recommendation systems (Wu et al., 2022; Yin et al., 2019) are prime examples.
-These systems evolve dynamically over time, exhibiting different characteristics at various temporal moments.
-Recently, there has been a marked increase in the representation learning tasks on temporal graphs in the graph research community.
-At the same time, a prominent graph learning tool, Graph Neural Network (GNN) (Scarselli et al., 2008), has been developed to model node, link, and graph tasks.
-GNNs generally learn node embeddings by iteratively aggregating embeddings from neighboring nodes.
-They have demonstrated exceptional performance in graph representation learning, achieving state-of-the-art results across numerous tasks.
-There remains a significant gap between static graphs and the increasingly prevalent temporal graphs.
-Temporal graphs incorporate discrete or continuous timestamps attached to edges, providing a more precise depiction of the graph evolution process.
-Meanwhile, many methodologies for static graphs are not applicable due to the additional constraints imposed by timestamps on causality.
-One can only utilize nodes and edges that precede a given time, implicitly resulting in numerous graph instances to analyze.
-Building on the success of time sequence modeling, Kumar et al.
-(2019); Trivedi et al.
-(2019); Rossi et al.
-(2020) propose memory-based temporal graph networks aimed at learning both short- and long-term dependencies.
-These methods, particularly Temporal Graph Networks (TGN), have achieved notable success in temporal node classification.
-Additionally, Transformer-based models, such as those proposed by Wang et al.
-(2021a); Xu et al.
-(2020), employ the multi-head attention mechanism to capture both cross and internal relationships within a temporal graph.
+Temporal graphs are increasingly employed in contemporary real-world applications like social and transaction networks, which evolve dynamically and exhibit distinct characteristics over time.
+Concurrently, Graph Neural Networks (GNNs) [scarselli2008graph] have emerged as prominent tools for graph representation learning, typically learning node embeddings by iteratively aggregating information from neighbors and demonstrating strong performance on various tasks.
+However, the temporal dimension introduces significant challenges; the discrete or continuous timestamps associated with graph edges define an evolutionary process and impose causality constraints, rendering many static GNN methodologies not directly applicable.
+This has led to the development of specialized temporal GNNs.
+Consequently, specialized temporal GNNs have been developed, often focusing on achieving both effective representation learning and computational efficiency for these large, dynamic structures.
+For instance, memory-based approaches like Temporal Graph Networks (TGN) [kumar2019predicting, trivedi2019dyrep, rossi2020temporal] are designed to efficiently learn short- and long-term dependencies, while Transformer-based models [wang2021tcl, xu2020inductive] utilize attention mechanisms to capture complex relationships.
 (a) a temporal graph
@@
 Thus with the same computation graph, we only need to utilize the extra node $b$’s embedding to distinguish $(u,w)$ and $(v,w)$.
-However, these models may exhibit inherent flaws when addressing link prediction tasks in graphs.
-Such tasks require the model to predict the existence of a target link given specific prerequisite information.
-The aforementioned approaches primarily focus on node-wise representation learning, using the node embeddings of source and destination nodes to predict the existence of a target link.
-While straightforward, this method often fails to precisely capture the relationship between two nodes and other more complex graph structure features.
-For example, in Figure 1, the node-wise method might struggle to predict whether node $w$ prefers to interact with node $u$ or $v$.
-Generally speaking, these models are constrained by their focus on node-level representation, limiting their ability to capture broader contexts and higher-order patterns.
-Considering these shortcomings, Zhang and Chen (2018) highlight the importance of pair-wise representation and propose labeling trick to mark the source and destination nodes when learning their pair-wise representation by GNN.
-Labeling trick (Zhang et al., 2021) is shown to greatly enhance GNN performance on static graph link prediction.
-Building on this, Wang et al.
-(2021c); Luo and Li (2022); Yu et al.
-(2023) develop models that leverage information from temporal surrounding nodes, successfully extending the pair-wise learning approach to dynamic graphs.
-They extract features using decreasing-timestamp random walks or joint neighborhood structures to generate multi-level embeddings for the center nodes, facilitating node classification and link prediction with pair-wise representations.
-However, these graph-based models often incur significant computational and memory costs due to extracting temporal neighborhoods and applying message passing on them for each node/link to predict, hindering their application in real-world, large-scale scenarios.
-In general, graph-based representation learning methods can significantly enhance model capabilities, yet their high computational cost limits widespread application.
-For example, the Temporal Graph Benchmark (TGB) (Huang et al., 2023) contains many high-quality real-world temporal graphs.
-However, most graph-based models choose to only evaluate on a subset of small graphs due to their unaffordable cost of training and evaluation on large-scale datasets.
-On the contrary, memory-based models (Rossi et al., 2020) update node representations sequentially following the event stream, thus is significantly faster than graph-based methods yet might lose important graph information.
-Motivated by these observations, we propose an expressive but efficient model, Temporal Neural Common Neighbor (TNCN), to combine the merits of both.
-By integrating a memory-based backbone with Neural Common Neighbor (Wang et al., 2023), TNCN learns expressive pairwise representations while maintaining high efficiency akin to memory-based models.
-Consequently, TNCN is suitable for large-scale temporal graph link prediction.
-We conducted experiments on five large-scale real-world temporal graph datasets from TGB.
-TNCN achieved new state-of-the-art results on three datasets and overall ranked first compared to 10 competitive baselines, demonstrating its effectiveness.
-To examine its scalability, we selected datasets with temporal edges ranging from $\mathcal{O}(10^{5})$ to $\mathcal{O}(10^{7})$ and node numbers ranging from thousands to millions.
-We found that TNCN achieved 1.9$\sim$4.7x speedup in training and 2.1$\sim$6.4x speedup in inference compared to graph-based models, while maintaining a similar scale of time consumption as memory-based models.
-2 Related Work
-2.1 Memory-based Temporal Graph Representation Learning
-Temporal graph learning has garnered significant attention in recent years.
-A classic approach in this domain involves learning node memory using continuous events with non-decreasing timestamps.
-Kumar et al.
-(2019) propose a coupled recurrent neural network model named JODIE that learns the embedding trajectories of users and items.
-Another contemporary work DyRep (Trivedi et al., 2019) aims to efficiently produce low-dimensional node embeddings to capture the communication and association in dynamic graphs.
-Rossi et al.
-(2020) introduces a memory-based temporal neural network known as TGN, which incorporates a memory module to store temporal node representations updated with messages generated from the given event stream.
-Apan (Wang et al., 2021b) advances the methodology by integrating asynchronous propagation techniques, markedly increasing the efficiency of handling large-scale graph queries.
-EDGE (Chen et al., 2021) emerges as a computational framework focusing on increasing the parallelizability by dividing some intermediate nodes in long streams each into two independent nodes while adding back their dependency by training loss.
-Chen et al.
-(2023) extend the update method for the node memory module, introducing an additional hidden state to record previous changes in neighbors.
-Complementing these efforts, additional contributions such as Edgebank (Poursafaei et al., 2022) and DistTGL (Zhou et al., 2023) have been directed towards formalizing and accelerating memory-based temporal graph learning methods.
-2.2 Graph-based Temporal Graph Representation Learning
-Subsequent works have incorporated the temporal neighborhood structure into temporal graph learning.
-CAWN (Wang et al., 2021c) employs random anonymous walks to model the neighborhood structure.
-TCL (Wang et al., 2021a) samples a temporal dependency interaction graph that contains a sequence of temporally cascaded chronological interactions.
-TGAT (Xu et al., 2020) considers the temporal neighborhood and feeds the features into a temporal graph attention layer utilizing a masked self-attention mechanism.
-NAT (Luo and Li, 2022) constructs a multi-hop neighboring node dictionary to extract joint neighborhood features and uses a recurrent neural network (RNN) to recursively update the central node’s embedding.
-This information is then processed by a neural network-based encoder to predict the target link.
-DyGFormer (Yu et al., 2023), instead, leverages one-hop neighbor embeddings and the co-occurrence of neighbors to generate features, which are well-patched and subsequently fed into a Transformer (Vaswani et al., 2017) decoder to obtain the final prediction.
-2.3 Link Prediction Methods
-Link prediction is a fundamental task in graph analysis, aiming to determine the likelihood of a connection between two nodes.
-Early investigations posited that nodes with greater similarity tend to be connected, which led to a series of heuristic algorithms such as Common Neighbors, Katz Index, and PageRank (Newman, 2001; Katz, 1953; Page et al., 1999).
-With the advent of GNNs, numerous methods have attempted to utilize vanilla GNNs for enhancing link prediction, revealing sub-optimal performance due to the inability to capture important pair-wise patterns such as common neighbors (Zhang and Chen, 2018; Zhang et al., 2021; Liang et al., 2022).
-Subsequent research has focused on infusing various forms of inductive biases to retrieve intricate pair-wise relationships.
-For instance, SEAL (Zhang and Chen, 2018), Neo-GNN (Yun et al., 2021), and NCN (Wang et al., 2023) have integrated neighbor-overlapping information into their design.
-BUDDY (Chamberlain et al., 2022) and NBFNet (Zhu et al., 2021) have concentrated on extracting higher-order structural information.
-Additionally, Mao et al.
-(2023); Li et al.
-(2024) have contributed to a more unified framework encompassing different heuristics.
+Despite their advancements, many specialized temporal GNNs primarily generate node-wise representations.
+While effective for certain tasks, this focus can be insufficient for link prediction, where accurately capturing the relationship between pairs of nodes is critical.
+Node-wise methods may fail to distinguish between nodes that have similar individual features or local neighborhoods but different propensities to connect with a target node (as illustrated in Figure 1), thereby overlooking crucial relational patterns.
+Acknowledging such limitations, approaches from static graph link prediction, like the labeling trick [zhang2018link, zhang2021labeling] that emphasizes pair-wise representations, have shown considerable success.
+Consequently, researchers have extended these pair-wise learning concepts to temporal graphs, often by leveraging information from the evolving local neighborhoods of node pairs [wang2021inductive, luo2022neighborhood, yu2023towards].
+However, these more expressive graph-based temporal models frequently incur substantial computational and memory costs, arising from the need to extract and process temporal neighborhood information for each prediction, which can hinder their application to large-scale scenarios.
+This trade-off between computationally demanding expressive models and more efficient but potentially less informative ones highlights a critical need.
+Addressing this, we propose the Temporal Neural Common Neighbor (TNCN) model, which is designed to achieve both high efficiency and strong expressive power for temporal link prediction.
+TNCN builds upon a memory-based backbone, ensuring operational efficiency comparable to sequential update models.
+Crucially, it incorporates a Neural Common Neighbor component [wang2023neural].
+This component is augmented with advanced operational techniques and extended for multi-hop common neighbor consideration, allowing TNCN to effectively model sophisticated link heuristics and learn detailed pairwise representations while retaining the efficiency of its memory-based foundation.
+As a result, TNCN is well-suited for large-scale temporal graph link prediction.
+[…]
 2 Preliminaries
 Definition 2.1.
-(Temporal Graph) Temporal graph can be typically categorized into two kinds, discrete-time (DTDG) and continuous-time (CTDG) dynamic graph.
-While DTDG can be represented as a special sequence of graph snapshots in the form of $\mathcal{G}=\{(\mathcal{G}_{1},t_{1}),(\mathcal{G}_{2},t_{2}),\cdots,(\mathcal{G}_{N},t_{N})\}$, it can always be transformed into its corresponding form in CTDG.
-So here we mainly focus on continuous time temporal graph.
-We usually represent a CTDG as a sequence of interaction events: $\mathcal{G}=\{(u_{1},v_{1},t_{1}),(u_{2},v_{2},t_{2}),\cdots,(u_{n},v_{n},t_{n})\}$, where $u,v$ stand for source and destination nodes and $\{t_{i}\}$ are chronologically non-decreasing.
-We use $\mathcal{V}$ to denote the entire node set, and $\mathcal{E}$ the entire edge set.
+(Temporal Graph) We mainly focus on the continuous time dynamic graph (CTDG).
+A CTDG can be typically represented as a sequence of interaction events: $\mathcal{G}=\{(u_{1},v_{1},t_{1}),\cdots,\\ (u_{n},v_{n},t_{n})\}$, where $u,v$ stand for source and destination nodes and $\{t_{i}\}$ are chronologically non-decreasing timestamps.
 Note that each node or edge can be attributed, that is, there may be node feature $x_{u}$ for $u$ or edge feature $e_{u,v}^{t}$ attached to the event $(u,v,t)$.
@@
 Definition 2.3.
-(Temporal Neighborhood) Given the center node $u$, the $k$-hop temporal neighbor set $(k\geq 0)$ before time $t$ is defined as $N_{k}^{t}(u)=\{v\ |\ \text{SPD}^{t}(u,v)=k\}$, where $\text{SPD}^{t}(u,v)$ refers to the shortest path distance using edges before time $t$.
-Then the $k$-hop common neighbors between node $u$ and $v$ are defined as $\text{CN}_{k}^{t}(u,v)=\mathop{\cup}\limits_{2(k-1)<k_{1}+k_{2}\leq 2k,k_{1},k_{2}\leq k}(N_{k_{1}}^{t}(u)\cap N_{k_{2}}^{t}(v))$, i.e., all nodes on at least one $2k$-distance path between $u,v$.
-Note that for $k=0$, we define the 00-hop common neighbor of $u$ and $v$ as $\text{CN}_{0}^{t}(u,v)=\{u,v\}$ if there exists at least one edge between $u$ and $v$ before time $t$, and $\emptyset$ otherwise.
-Finally, the $K$-hop temporal neighborhood is defined as $\mathop{\cup}\limits_{k=0}^{K}N_{k}^{t}(u)$.
+(Temporal Neighborhood) Given the center node $u$, the $k$-hop temporal neighbor set $(k\geq 0)$ before time $t$ is defined as $N_{k}^{t}(u)$.
+A node $v$ is in $N_{k}^{t}(u)$ if there exists a $k$-length path between $u$ and $v$, i.e.
+$\exists(u,w_{1},w_{2},\cdots,w_{k-1},v)$ where $w_{i}\neq w_{j},\forall i\neq j$.
+We also define the $(\mathbf{i},\mathbf{j})$-hop common neighbor set as follows: $w$ is an $(i,j)$-hop temporal common neighbor of $u$ and $v$ at time $t$ if $w\in N_{i}^{t}(u)$ and $w\in N_{j}^{t}(v)$.
+For simplicity we will denote the set as $\text{CN}_{(i,j)}^{t}(u,v)=N_{i}^{t}(u)\cap N_{j}^{t}(v)$.
+Note that for $i=0$ (or $j=0$ similarly), we define the 0-hop temporal neighbor set as $N_{0}^{t}(u)=\{u\}$, and the $(0,j)$-hop common neighbor of $u$ and $v$ as $\text{CN}_{(0,j)}^{t}(u,v)=N_{0}^{t}(u)\cap N_{j}^{t}(v)=\{u\}\cap N_{j}^{t}(v)$.
+Finally, the $K$-hop temporal neighborhood of node $u$ at time $t$ is defined as: $\mathop{\cup}\limits_{k=0}^{K}N_{k}^{t}(u)$.
+With $(i,j)$-hop neighborhood information, we can perceive the local structure to a large extent and distinguish the difference between multi-hop common neighbors more precisely.
+Definition 2.4.
+Memory-based Backbone.
+Memory-based backbone has been widely adopted by various methods like [kumar2019predicting, rossi2020temporal] to tackle dynamic graph learning.
+Its core component is the memory module that stores the node memory representations up to a certain time $t$.
+When a new event occurs, the memory of the source and destination nodes is updated with the message produced by the event.
+The computation can generally be represented as follows:
+$msg_{src}^{t}(u,v)=\textit{msgfunc}_{src}(e^{t}_{u,v}),$ $mem^{t}_{u}=\textit{upd}_{src}(mem^{t-}_{u},msg_{src}^{t}(u,v));$ (1)
+$msg_{dst}^{t}(u,v)=\textit{msgfunc}_{dst}(e^{t}_{u,v}),$ $mem^{t}_{v}=\textit{upd}_{dst}(mem^{t-}_{v},msg_{dst}^{t}(u,v)).$
+where $msg^{t}$ stands for the message of the event, $mem^{t-}_{u}$ for the embedding of node $u$ before time $t$.
 3 Methodology
-In this section, we introduce our Temporal Neural Common Neighbor (TNCN) model.
-TNCN comprises several key modules: the Memory Module, the Temporal CN Extractor, and the NCN-based Prediction Head.
-Special attention is given to the Temporal CN Extractor, designed to efficiently extract temporal neighboring structures and obtain multi-hop common neighbor information.
-The pipeline of TNCN is illustrated in Figure 2.
+Now we introduce our Temporal Neural Common Neighbor model.
+TNCN comprises several key modules: the classic Memory Module, the Temporal NCN Module with efficient CN Extractor, and the NCN-based Prediction Head.
+Special attention will be given to the Temporal CN Extractor, which is designed to efficiently extract temporal neighboring structures and obtain multi-hop CN information.
+The pipeline is illustrated in Figure 2.
+A pseudocode is attached in Appendix F.
 Figure 2: Pipeline of TNCN.
 TNCN operates through a sequential update and prediction framework that processes successive batches of messages.
-During the update phase, TNCN updates the neighbor dictionary and the node memory representations.
-In the prediction phase, the model retrieves neighbors to identify common neighbors, thereafter leveraging the representations of the target nodes and their common neighbors for prediction.
+During the update phase, TNCN updates the neighbor dictionary and the node memory.
+In the prediction phase, the model retrieves neighbors to identify common neighbors, leveraging the representations of the target nodes and their CNs for prediction.
 3.1 Memory Module
-The memory module stores node memory representations up to time $t$.
-When a new event occurs, the memory of the source and destination nodes is updated with the message produced by the event.
-The computation can generally be represented by the following formulas:
-$msg_{src}^{t}(u,v)=\textit{msgfunc}_{src}(e^{t}_{u,v}),$ $msg_{dst}^{t}(u,v)=\textit{msgfunc}_{dst}(e^{t}_{u,v}).$ (1)
-$mem^{t}_{u}=\textit{upd}_{src}(mem^{t-}_{u},msg_{src}^{t}(u,v)),$ $mem^{t}_{v}=\textit{upd}_{dst}(mem^{t-}_{v},msg_{dst}^{t}(u,v)),$
-where $mem^{t-}_{u}$ stands for the embedding of node $u$ before time $t$.
-Here, msgfunc is a learnable function, such as a linear projection or a simple identity function.
-Note that for different edge directions, i.e., from source to destination and vice versa, the msgfunc and updfunc can be learned separately.
-The memory module aids the model in managing both long-term and short-term dependencies, thereby reducing the likelihood of forgetting.
-During training and inference, node memory evolves dynamically as events occur.
-Updates to this module reflect the dynamic nature of the temporal graph.
-4.2 Temporal CN Extractor
-Our Temporal CN Extractor can efficiently perform multi-hop common neighbor extraction and aggregate their neural embeddings.
+Different from the static GNN used in traditional NCN, our model TNCN adopts a memory-based backbone from [rossi2020temporal] to efficiently store and update the node memory, eliminating the need for repeated computation of node embeddings within successive temporal batches.
+Conforming to the standard pipeline in section 2.4 of processing the node memory, for later link prediction or other downstream tasks, node embeddings can be obtained from their memory:
+$emb^{t}_{u}=NN(mem^{t-}_{u},\mathop{\cup}\limits_{v\in N^{t}_{1}(u)}mem^{t-}_{v},\mathop{\cup}\limits_{t^{\prime}<t}[e^{t^{\prime}}_{u,v}\ ||\ \phi(t-t^{\prime})]),$ (2)
+where $||$ stands for the concatenation operation.
+Here NN has multiple choices, like Identity or simple static GNN [defferrard2016convolutional, velickovic2017graph, hamilton2017inductive].
+In our implementation, we adopt Graph Transformer Convolution [shi2020masked], which can pay more attention to the relation between different nodes.
+And we choose the time encoding $\phi$ presented in Time2Vec [kazemi2019time2vec] as TGN does.
+3.2 Temporal NCN Module with Efficient CN Extractor
+Our Temporal NCN Module can efficiently perform multi-hop common neighbor extraction with its CN extractor, and aggregate their neural embeddings to attain the node features.
 Extended Common Neighbor.
-The definition of multi-hop common neighbors (CN) is given in Definition 3.3, extending the traditional 1-hop CN (i.e., nodes on 2-paths between $u$ and $v$) to $k$-hop CN (nodes on $2k$-paths between $u$ and $v$).
-Additionally, we define zero-hop CN to be $u$ and $v$ themselves, i.e., if there is at least one previous event between node $u$ and node $v$, then $u$ and $v$ are considered each other’s 0-hop common neighbors.
-The 0-hop CN not only records the historical interactions between two nodes, but also reveals the frequency of their interactions.
+The definition of multi-hop common neighbors (CN) is given in Definition 2.3, extending the traditional (1,1)-hop CN (i.e., nodes on 2-paths between $u$ and $v$) to arbitrary $(i,j)$-hop CN.
+Additionally, we define the zero-hop neighbor of a central node, i.e., $u$ is considered as a neighbor of itself, which will be utilized to calculate CNs with other nodes.
+Given source node $u$ and target node $v$, the (0,1)-hop and (1,0)-hop CN not only records the historical interactions between two nodes, but also reveals the frequency of their interactions.
 Efficient CN Extractor.
-The CN Extractor is a crucial component of the TNCN model, contributing significantly to its high performance.
+The CN Extractor is a crucial component of the TNCN model, contributing significantly to its high performance and scalability.
 It can efficiently gather pertinent information about a given center node and extract multi-hop common neighbors for a source-destination pair.
@@
 This allows us to maintain a record of all historical interactions up to a certain timestamp, effectively constructing a dynamic lookup dictionary for fast retrieval during subsequent inference.
-To conserve memory, we can opt to save only the most recent events and relevant nodes.
-During the inference stage, TNCN extracts temporal neighborhood information, i.e., historical interactions related to the query node batch.
-This information is organized as a Sparse Tensor, on which we perform self-multiplication to generate high-order adjacency connectivity.
-We use sparse tensor operations to obtain multi-hop common neighbors (CNs) and their embeddings, separating CNs from different hops.
-To extract the 0-hop common neighbors mentioned above in a unified way, we introduce an additional self-loop for each related node.
-This operation allows us to use the same tensor multiplications to obtain 0-hop CNs.
+To strike a balance between memory consumption and model capacity, we save only the most recent $K$ events and relevant nodes for each center node, where $K$ is a hyperparameter determined by the specific dataset.
+To implement an efficient batch CN extractor, we organize the historical interactions in a Sparse Tensor, representing the temporal adjacency matrix.
+Then we perform self-multiplication to generate high-order adjacency connectivity.
+Sparse matrix hadamard product is finally employed to obtain separate $(i,j)$-hop CNs.
+All these operations can be efficiently implemented by sparse tensor operators and are supported by GPU to facilitate fast, batch processing.
+Now we show the detailed procedure as follows.
+For some special and higher-order cases analysis, please refer to Appendix G.
+Details of Common Neighbor Extraction.
+Our temporal CN extractor begins with a sparse matrix $A$ constructed from the interactions of related nodes.
+We then include three stages to precisely generate arbitrary $(i,j)$-hop CNs:
+1.
+Generate up to $k$-hop neighbors.
+The original matrix $A$ only includes 1-hop neighbors.
+To extend this, we: (a) Use self-loops for 0-hop neighbors, denoted as $A^{0}$.
+(b) Perform sparse matrix multiplication (i.e., $A^{k}$) to include arbitrary k-hop neighbors.
+Combining them, we obtain an updated neighborhood matrix set $\hat{A}=\{A^{i}\}_{i=0}^{K}$.
+2.
+Extract neighbors for each source and destination node with corresponding indices in the same batch.
+Assume that we require the $k$-th hop neighbors of node $u$, then vector $A^{k}[id(u)]$ is the result, where $id(u)$ stands for the reindexed id for node $u$.
+$A^{k}[id(u)][id(v)]=w>0$ if $v$ is a $k$-hop neighbor of $u$, otherwise this element is 0.
+$w$ represents the historical interaction frequency.
+3.
+Obtain arbitrary $(i,j)$-hop CNs.
+We can perform hadamard product of $A^{i}[id(u)]$ and $A^{j}[id(v)]$ to acquire different hops of CNs.
+The operator can extract corresponding CNs for source-destination node pairs in a batch parallelly.
+By re-indexing the node IDs when generating $\hat{A}$ to prevent conflicts, the CN extractor can conduct the sparse matrix calculation, which are all performed in a Torch style that supports batch operations, thus enhancing parallelism and efficiency.
 The utilization of Multi-hop Common Neighbors significantly boosts TNCN’s performance, resulting in higher scores in temporal link prediction tasks.
 Furthermore, by employing sparse tensors, our model achieves substantial reductions in both storage requirements and computational complexity, thereby decreasing time consumption and enhancing efficiency.
+We also give a comparison between our TNCN and traditional NCN in Table 12 in Appendix D.4.
 3.3 NCN-based Prediction Head
-For source and destination nodes, we perform an element-wise product:
-$X^{t}_{u,v}=emb_{u}^{t}\otimes emb_{v}^{t}.$ (2)
-For multi-hop CN nodes, we aggregate their embeddings in each hop with sum pooling:
-$NCN_{k}(u,v)=\mathop{\oplus}\limits_{w\in\text{CN}^{t}_{k}(u,v)}emb_{w}^{t}.$ (3)
+We finally construct our NCN-based representation as follows.
+For source and destination nodes, we perform an element-wise product.
+For multi-hop CN nodes, we aggregate their embeddings in each hop with sum pooling.
+$X^{t}_{u,v}=emb_{u}^{t}\otimes emb_{v}^{t},\quad NCN_{(i,j)}(u,v)=\mathop{\oplus}\limits_{w\in\text{CN}^{t}_{(i,j)}(u,v)}emb_{w}^{t}.$ (3)
 These embeddings are then concatenated as the final pair-wise representation:
-$repr(u,v)$ $=[X^{t}_{u,v}\ ||\ (\mathop{||}\limits_{k=0}^{K})NCN_{k}(u,v)].$ (4)
-In the above, $\otimes$, $\oplus$, and ||||| | stand for element-wise product, element-wise summation, and concatenation of vectors, respectively.
-The pair-wise representation $repr(u,v)$ for nodes $u$ and $v$ is fed to a projection head to output the final link prediction.
+$repr(u,v)$ $=[X^{t}_{u,v}\ ||\ (\mathop{||}\limits_{i,j}^{K})NCN_{(i,j)}(u,v)].$ (4)
+We have used $\otimes$, $\oplus$, and $||$ to denote element-wise product, element-wise summation, and concatenation of vectors, respectively.
+The pair-wise representation $repr(u,v)$ for nodes $u$ and $v$ will be fed to a projection head to output the final link prediction.
 4 Efficiency and Effectiveness of TNCN
@@
 These advantages are demonstrated through an analysis of two core components within the framework for temporal graph link prediction: graph representation learning and link prediction methods.
-We categorize graph representation learning modules into two types: memory-based and $k$-hop-subgraph-based, according to the temporal scope of evolved events.
-Memory-based modules exhibit superior time complexity while maintaining good expressiveness in some situations, striking a balance between efficiency and performance.
-Furthermore, we highlight the deficiencies of existing link prediction methods on temporal graph learning and introduce the extended common neighbor approach.
-This method serves as a complementary addition for learning pair-wise representations while eliminating the necessity for message passing on entire graphs.
-Both our graph representation learning and link prediction techniques are designed with a unified optimization objective: to avoid message passing on entire subgraphs in favor of non-repetitive operations, culminating in a cohesive solution that is both efficient and effective.
-5.1 Graph Representation Learning
-Graph representation learning aims to develop an embedding function, denoted as $Emb$, which learns an embedding for each node encoding its structural and feature information within the graph.
+Temporal graph representation learning aims to develop an embedding function, denoted as $Emb$, which learns an embedding for each node encoding its structural and feature information within the graph.
 Specifically, given a new event represented as $(u,v,t)$, the function $Emb$ leverages prior events to generate meaningful embeddings.
-Approaches in this domain diverge in their handling of temporal dynamics; some opt to maintain a dynamic embedding for each node that is incrementally updated with each new event, while others choose to recalculate node embeddings by considering the entire historical context of events, thereby providing a more comprehensive reflection of past interactions.
-We classify these methodologies into two distinct types based on their operational mechanisms.
+We first categorize graph representation learning approaches into two types: memory-based and $k$-hop-subgraph-based, according to their temporal scope of evolved events.
 Definition 4.1.
-Memory-based approach. Given a new event $(u,v,t)$, if $Emb$ conforms to the following form, the method is referred to as a memory-based approach.
-$Emb(u,t)=f_{\textit{emb}}(Mem(u,t^{\prime})),~{}$ $Mem(u,t)=f_{\textit{mem}}(Mem(u,t^{\prime}),Mem(v,t^{\prime}),e_{u,v}^{t},t-t^{\prime}),$ (5)
-$Emb(v,t)=f_{\textit{emb}}(Mem(v,t^{\prime})),~{}$ $Mem(v,t)=f_{\textit{mem}}(Mem(v,t^{\prime}),Mem(u,t^{\prime}),e_{v,u}^{t},t-t^{\prime}),$
-where $f_{\textit{emb}}$ and $f_{\textit{mem}}$ are two learnable functions.
+Memory-based approach.
+Given a new event $(u,v,t)$, if $Emb$ conforms to the following form, the method is referred to as a memory-based approach, which opts to maintain a dynamic, incrementally updated embedding for each node.
+$Emb(u,t)=f_{\textit{emb}}(Mem(u,t^{\prime})),~~~~Emb(v,t)=f_{\textit{emb}}(Mem(v,t^{\prime})),$ (5)
+where the $Mem$ can be obtained from the pipeline in Definition 2.4, and $f_{\textit{emb}}$ is a learnable function.
 Definition 4.2.
-$k$-hop-subgraph-based approach. Given a new event $(u,v,t)$, if $Emb$ conforms to the following form, the method is defined as a subgraph-based approach.
+$k$-hop-subgraph-based approach.
+Given a new event $(u,v,t)$, if $Emb$ conforms to the following form, the method is defined as a subgraph-based approach, which chooses to recalculate node embeddings by considering the entire historical events.
 $Emb(u,t)=f_{\textit{emb}}(\mathcal{G}_{u,<t}^{k}),~~~~Emb(v,t)=f_{\textit{emb}}(\mathcal{G}_{v,<t}^{k}),$ (6)
 where $\mathcal{G}_{u,<t}^{k}$ is a subgraph induced from $\mathcal{G}$ by node $u$’s $k$-hop temporal neighborhood $\mathop{\cup}\limits_{k=0}^{K}N_{k}^{t}(u)$, containing only the edges (events) with time $t^{\prime}<t$, and $f_{\textit{emb}}$ is a learnable function.
-Effectiveness.
-The analysis begins by assessing the effectiveness of the two paradigms. To do so, we first introduce the concept of $k$-hop event (LOV´ASZ et al., 1993).
+4.1 Effectiveness
+The analysis begins by assessing the effectiveness of the two paradigms. To do so, we first introduce the concept of $k$-hop event [LOVASZ1993RandomWO].
 Definition 4.3.
-($k$-hop event & monotone $k$-hop event) A $k$-hop event is a sequence of consecutive edges $\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$ connecting the initial node $u_{0}$ to the final node $u_{k}$.
+$k$-hop event & monotone $k$-hop event.
+A $k$-hop event is a sequence of consecutive edges $\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$ connecting the initial node $u_{0}$ to the final node $u_{k}$.
 For example, $\{(u,x,t^{\prime}),(x,v,t)\}$ is a $2$-hop event.
@@
   • $k$-hop-subgraph-based approaches can encode any $k^{\prime}$-hop events with $k^{\prime}\leq k$.
-From Theorem 5.4, we can conclude that 1) memory-based approaches have superior expressiveness in encoding $k$-hop events compared to 1-hop-subgraph-based approaches, 2) memory-based approaches have superior expressiveness in encoding monotone $k$-hop events than $k^{\prime}$-hop-subgraph-based approaches when $k^{\prime}<k$, and 3) $k$-hop-subgraph-based approaches are not less expressive than memory-based approaches when $k$ is large enough.
+The proof is in Appendix I.
+From Theorem 4.4, we can conclude that 1) memory-based approaches have superior expressiveness in encoding $k$-hop events compared to 1-hop-subgraph-based approaches, and 2) memory-based approaches have superior expressiveness in encoding monotone $k$-hop events than $k^{\prime}$-hop-subgraph-based approaches when $k^{\prime}<k$.
 While the memory-based approach does not consistently rival the expressiveness of the $k$-hop-subgraph paradigm, it possesses advantages in monotone events and long-history scenarios (where $k$-hop subgraphs would be unaffordable to extract).
-Efficiency.
+Corollary 4.5.
+If we use up to $k$ hop neighborhood information of central node $u$, then TNCN can capture at least $(k+1)$-hop subgraph information around $u$.
+This is because TNCN with memory-based backbone can obtain additional 1-hop information regardless of the time monotony, i.e.
+arbitrary central node can interact with its neighbor when the edge between them exists.
+This can extend TNCN’s capability for free.
+We also demonstrate the effectiveness of TNCN in the following theorem 4.6 with the proof in Appendix I.
+The first part shows that it can capture three important pairwise features commonly used as effective link prediction heuristics, namely Common Neighbors (CN), Resource Allocation (RA), and Adamic-Adar (AA) [newman2001clustering, adamic2003friends, zhou2009predicting].
+The experimental results in Appendix D further validate this claim.
+In the second part we reveal that TNCN is strictly more expressive than some traditional temporal graph networks including Jodie [kumar2019predicting], DyRep [trivedi2019dyrep], TGN [rossi2020temporal] and TGAT [xu2020inductive] under the same condition, which are widely used as baselines.
+Theorem 4.6.
+(Expressivity of TNCN)
+  1. 1.
+    TNCN is strictly more expressive than CN, RA, and AA.
+  2. 2.
+    TNCN is strictly more expressive than Jodie with the same dimension of time encoding, DyRep with the same aggregation function, TGAT with the same attention layers and neighbors, and TGN under identical condition for all module choices.
+From this theorem we can find that TNCN extends the previous generic memory-based framework of temporal graph networks with explicitly adopting the neural embeddings of common neighbors.
+This method can serve as a complementary addition for learning pair-wise representations.
+4.2 Efficiency
 We then turn our attention to the efficiency of the two approaches.
@@
 Conversely, in the $k$-hop-subgraph-based method, an event may be employed multiple times, as it is revisited in different nodes’ temporal neighborhood and repeated been processed within each subgraph’s encoding (such as message passing) process.
-This discrepancy leads to divergent cumulative frequencies of event utilization throughout the learning process and results in the huge efficiency advantage of memory-based methods.
-We formalize this observation in the following:
+This discrepancy leads to divergent cumulative frequencies of event utilization throughout the learning process, resulting in the huge efficiency advantage of memory-based methods.
+We formalize this observation as:
 Theorem 4.7.
@@
   • For $k$-hop-subgraph-based approaches with $k=2$, the upper-bound time complexity is $\mathcal{O}\left(\left(\frac{|\mathcal{E}|^{2}}{|\mathcal{N}|}+|\mathcal{E}||\mathcal{N}|\right)^{\frac{3}{2}}\right)$.
-Following Theorem 5.5, it becomes evident that the computational overhead incurred by a memory-based method is significantly lower than that of a $k$-hop-subgraph-based method, particularly as the value of $k$ increases.
+The proof is attached in Appendix I with part of the proof based on a classic conclusion from DECAEN1998245 in the graph theory.
+Following Theorem 4.7, it becomes evident that the computational overhead incurred by a memory-based method is significantly lower than that of a subgraph-based method, particularly as $k$ increases.
 These results highlight the advantages of memory-based methods in mitigating the computational efficiency challenges associated with large-scale temporal graphs.
-5.2 Link Prediction Technique
-With the node embeddings obtained from the graph representation learning step, link prediction techniques involve aggregating the node embeddings in some way into link representations for link prediction.
-Most previous methods simply concatenate the source and destination node embeddings as the link representation $Emb(u,t)\mathop{||}Emb(v,t)$, losing a great amount of pairwise features (as illustrated in Figure 1).
-In order to use labeling trick, an alternative approach involves extracting a separate subgraph for each link to predict, which leads to good results but also significantly increases the complexity.
-TNCN, on the contrary, utilizes neural common neighbors as a decoupled and flexible solution.
-This approach can be seamlessly integrated into existing memory-based methods with minimal computational overhead, while retaining important structural information for link prediction.
-Effectiveness.
-In the following, we first demonstrate the effectiveness of TNCN by showing that it can capture three important pairwise features commonly used as effective link prediction heuristics, namely Common Neighbors (CN), Resource Allocation (RA), and Adamic-Adar (AA) (Newman, 2001; Adamic and Adar, 2003; Zhou et al., 2009), borrowing from NCN (Wang et al., 2023).
-Theorem 5.6.
-TNCN is strictly more expressive than CN, RA, and AA.
-Essentially, TNCN uses node memory to substitute the constant or degree-based values in the three features, and the memory update scheme is sufficient to learn such values.
-In comparison, an approach that merely concatenates node-wise representations proves inadequate in capturing these heuristics.
-Efficiency.
-We then address the efficiency of TNCN.
-The computation of TNCN can be divided into two primary components: the generation of neighbor nodes’ embeddings and the execution of common neighbor lookups.
-Concerning the former, the memory-based approach intrinsically tracks all node embeddings, thus obviating the need for re-computation.
-As for the latter, we implement a fast common neighbor search algorithm by leveraging a sparse matrix structure that supports batch operations.
-Collectively, these factors contribute to the minimal additional overhead of TNCN compared to memory-based methods.
+So based on the memory backbone, our TNCN furthur utilizes an efficient CN extractor, eliminating the necessity for message passing on entire graphs.
+Consequently it achieves a unified optimization objective: to avoid message passing on entire subgraphs in favor of non-repetitive operations.
+To summarize, TNCN introduces the extended common neighbor approach with the efficient CN extractor.
+This method serves as a complementarity for learning pair-wise representations, while culminating in a cohesive solution that is both efficient and effective.
+5 Related Work
+Memory-based Temporal Graph Representation Learning.
+Memory-based models learn node memory using continuous events with non-decreasing timestamps.
+Researchers have proposed several memory-based methods including JODIE [kumar2019predicting], DyRep [trivedi2019dyrep] and TGN [rossi2020temporal].
+These methods are superior in higher efficiency, while lacking in capturing structural information.
+Graph-based Temporal Graph Representation Learning.
+Subsequent works have incorporated the neighborhood structure into temporal graph learning.
+CNE-N [cheng2024co] employs a hash table to map interactions and computes co-neighbor encodings to predict future links.
+Unlike our TNCN which directly utilizes CN embeddings in prediction, CNE-N simply counts neighbor numbers.
+Additionally, while CNE-N manages recent interactions through hash tables, TNCN adopts a monotonic storage scheme.
+NAT [luo2022neighborhood] similarly builds a multi-hop node dictionary to compress neighbors, but it may suffer from hash collisions.
+DyGFormer [yu2023towards] encodes one-hop neighbors and their co-occurrences, then uses a Transformer for predictions, requiring repeated neighborhood sampling and computations.
+It only models 1-hop neighbor frequency and does not use CN embeddings.
+In contrast, TNCN supports multi-hop neighbors and includes their embeddings, enabling richer representations.
+We also include detailed introduction to more related works in Appendix B.
 6 Experiments
@@
 6.1 Experimental Settings
-Datasets
-We evaluate our model on five large-scale real-world datasets for temporal link prediction from the Temporal Graph Benchmark (Huang et al., 2023).
+Datasets.
+We evaluate our model on five large-scale real-world datasets for temporal link prediction from the Temporal Graph Benchmark [huang2023temporal].
 These datasets span several distinct fields: co-editing network on Wikipedia, Amazon product review network, cryptocurrency transactions, directed reply network of Reddit, and crowdsourced international flight network.
@@
 We set the evaluation metric as Mean Reciprocal Rank (MRR) consistent with the TGB official leaderboard.
-Baselines
-We systematically evaluate our proposed model against a diverse set of baselines known for their strong capacity to represent temporal graph dynamics.
-These include: a heuristic algorithm Edgebank (Yu et al., 2023), memory-based models DyRep (Trivedi et al., 2019) and TGN (Rossi et al., 2020) that obviate the need for frequent temporal subgraph sampling, and GraphMixer (Cong et al., 2023) which primarily employs an MLP-mixer.
-Additionally, we include various graph-based models such as CAWN (Wang et al., 2021c), NAT (Luo and Li, 2022), DyGFormer (Yu et al., 2023), TGAT (Xu et al., 2020), and TCL (Wang et al., 2021a), which learn from neighborhood structure information.
-To reconcile the trade-off between model efficacy and computational overhead, we evaluate two TNCN variants: the “TNCN-standard” (TNCN(S)), which incorporates up to 1-hop common neighbor information, and the more complex “TNCN-large” (TNCN(L)), which uses up to 2-hop common neighbors.
+Baselines.
+We systematically evaluate our proposed model TNCN against a diverse set of baselines: a heuristic algorithm Edgebank [yu2023towards], memory-based models Jodie [kumar2019predicting], DyRep [trivedi2019dyrep] and TGN [rossi2020temporal] that obviate the need for frequent temporal subgraph sampling, and GraphMixer [cong2023we] which employs an MLP-mixer.
+We also include various graph-based models such as CAWN [wang2021inductive], TGAT [xu2020inductive], TCL [wang2021tcl], NAT [luo2022neighborhood], DyGFormer [yu2023towards], CNE-N [cheng2024co] and TPNet [lu2024improvingtemporallinkprediction], which learn from neighborhood structure information.
+Here we evaluate our TNCN under two similar but different settings, the official setting (“official”) and the new setting (“ns”).
+“*-official” strictly complies to the official setting of TGB evaluation policy, using both streaming setting and lag-one scheme for both memory update and neighborhood awareness.
+Streaming setting means the information of the validation and test sets can only be employed for updating the memory without any back propagation.
+Lag-one scheme implies that the model can access only the information from before the current batch for predictions; in other words, the latest usable batch is the previous one.
+This applies to not only the memory, but also the neighborhood awareness.
+“*-ns” obeys the streaming setting but considers the interactions within the same batch before the current prediction time.
+This allows the model to utilize more recent neighborhood information, potentially giving it unfair advantages in datasets where recent interactions are crucial.
+Methods “*-official” use the former setting while others report the latter.
+For a fair evaluation and comparison, here we display the performance of our TNCN under both settings.
+Table 1: Test Performance of different models under MRR metric.
+The top three are emphasized by red, blue and bold fonts.
+‘-’ denotes scenarios where a specific method was either not applied to the dataset or was unable to complete the validation and testing phases within a reasonable timeframe.
+Model Wiki Review Coin Comment Flight
+JODIE 0.631 $\pm$ 1.69 0.414 $\pm$ 0.15 - - -
+DyRep 0.519 $\pm$ 1.95 0.401 $\pm$ 0.59 0.452 $\pm$ 4.60 0.289 $\pm$ 3.30 0.556 $\pm$ 1.40
+TGAT 0.599 $\pm$ 1.63 0.196 $\pm$ 0.23 0.609 $\pm$ 0.57 0.562 $\pm$ 2.11 -
+TGN-official 0.528 $\pm$ 0.06 0.387 $\pm$ 0.02 0.737 $\pm$ 0.03 0.622 $\pm$ 0.02 0.705 $\pm$ 0.02
+TGN-ns 0.689 $\pm$ 0.53 0.375 $\pm$ 0.23 0.586 $\pm$ 3.70 0.379 $\pm$ 2.10 -
+CAWN 0.730 $\pm$ 0.60 0.193 $\pm$ 0.10 - - -
+EdgeBank(tw) 0.633 0.029 0.574 0.149 0.387
+EdgeBank(un) 0.525 0.023 0.359 0.129 0.167
+TCL 0.781 $\pm$ 0.20 0.165 $\pm$ 1.85 0.687 $\pm$ 0.30 0.701 $\pm$ 0.83 -
+GraphMixer 0.598 $\pm$ 0.39 0.369 $\pm$ 1.50 0.756 $\pm$ 0.27 0.762 $\pm$ 0.17 -
+NAT 0.749 $\pm$ 1.00 0.341 $\pm$ 2.00 - - -
+DyGFormer 0.798 $\pm$ 0.42 0.224 $\pm$ 1.52 0.752 $\pm$ 0.38 0.670 $\pm$ 0.14 -
+CNE-N 0.802 $\pm$ 0.20 0.261 $\pm$ 0.25 0.772 $\pm$ 0.21 0.790 $\pm$ 0.14 -
+TPNet 0.827 $\pm$ 0.01 - 0.832 $\pm$ 0.01 0.825 $\pm$ 0.06 0.884 $\pm$ 0.01
+TNCN-official 0.724 $\pm$ 0.01 0.419 $\pm$ 0.09 0.770 $\pm$ 0.06 0.727 $\pm$ 0.12 0.817 $\pm$ 0.04
+TNCN-ns 0.803 $\pm$ 0.01 0.427 $\pm$ 0.06 0.771 $\pm$ 0.04 0.705 $\pm$ 0.12 0.831 $\pm$ 0.03
 6.2 Experimental Results
-Table 1: Test Performance of different models under MRR metric.
-The top three are emphasized by red, blue and black fonts.
-‘NA’ denotes scenarios where a specific method was either not applied to the dataset or was unable to complete the validation and testing phases within a reasonable timeframe.
-Model Wiki Review Coin Comment Flight
-TGN 0.528 0.387 0.737 0.622 0.705
-DyGformer 0.798 0.224 0.752 0.670 NA
-NAT 0.749 0.341 NA NA NA
-CAWN 0.711 0.193 NA NA NA
-Graphmixer 0.118 0.521 NA NA NA
-TGAT 0.141 0.355 NA NA NA
-TCL 0.207 0.193 NA NA NA
-DyRep 0.050 0.220 0.452 0.289 0.556
-Edgebank(tw) 0.571 0.025 0.580 0.149 0.387
-Edgebank(un) 0.495 0.023 0.359 0.129 0.167
-TNCN-standard 0.720 0.298 0.739 0.692 0.803
-TNCN-large 0.717 0.317 0.764 0.662 NA
 Reply to Q1: TNCN possesses remarkable performance.
-We conducted comprehensive evaluations of prevailing methods on the TGB.
+We conducted comprehensive evaluations of prevailing methods on TGB.
 The main results are summarized in Table 1.
-It is evident from the table that TNCN attains new SOTA performance on three out of five datasets.
-Additionally, TNCN demonstrates competitive results on the remaining two datasets.
-TNCN almost consistently surpasses memory-based models such as TGN and DyRep, which can be attributed to its utilization of supplementary structural information.
-Even in comparison to powerful graph-based models, including NAT and DyGFormer, TNCN still matches or exceeds their performance, underscoring the effectiveness of its integration of node memory and pair-wise features.
-The only dataset where TNCN exhibits relatively poorer performance is the Review dataset.
-This can be ascribed to the dataset’s predominant reliance on feature embedding and its high "surprise" metric, wherein prior events have a diminished correlation with subsequent events, potentially reducing the impact of CNs.
+It is evident from the table that TNCN attains new SOTA performance on Review dataset.
+Additionally, TNCN demonstrates competitive results on the remaining datasets, ranking 2nd or/and 3rd on Wiki, Coin and Flight.
+TNCN almost consistently surpasses both memory-based models such as TGN and DyRep, and graph-based methods like DyGFormer and NAT.
+The dataset where TNCN still exhibits a large gap from the top-three baselines is Comment.
+This may be ascribed to the high “surprise index” of this non-bipartite dataset, wherein prior events have a diminished correlation with subsequent events, potentially reducing the impact of CNs.
+(More details about Surprise index are in Appendix A.)
+To obtain a more comprehensive evaluation, we have also conducted experiments on the previous small and medium datasets under both transductive and inductive settings.
+In these experiments, TNCN achieves 6 new SOTA out of 7 datasets in transductive setting and 3 in inductive setting, further facilitating its strong performance.
+The overall performance and some additional results (such as TGN with heuristics, etc.) can be referred to Appendix D.
 Reply to Q2: TNCN shows great scalability on large datasets.
+To evaluate computational efficiency, we collected the time consumption on Wiki and Review datasets, as depicted in Figure 3.
+Compared with memory-based methods, TNCN exhibits a comparable order of magnitude in terms of time consumption.
+However, when benchmarked against graph-based models, TNCN demonstrates a substantial acceleration, achieving approximately 2.5 to 5.9 times speedup during the training phase and a 1.8 to 30.3 times increase in inference speed.
+In the Table  13 we provide a full comparison of the time consumption for different models over the five TGB datasets.
+Notably, the scalability concerns become even more evident as the size of the dataset expands; several graph-based models cannot complete the validation and testing processes within a reasonable time budget.
+The primary factors contributing to TNCN’s efficiency are the synergistic, time-efficient design of its two core components and the implementation of the Efficient CN Extractor that facilitates batch operations through parallel processing.
+For detailed statistics about TNCN and NAT, please refer to Appendix E.1.
 (a) tgbl-wiki (log scale)
 (b) tgbl-review (linear scale)
-Figure 3: Time Consumption of Memory and Graph-based Method on Wiki and Review Datasets
-To evaluate computational efficiency, we collected the time consumption when applying all baselines to the Wiki and Review datasets, as depicted in Figure 3.
-Compared with memory-based methods, TNCN exhibits a comparable order of magnitude in terms of time consumption.
-However, when benchmarked against graph-based models, TNCN demonstrates a substantial acceleration, achieving approximately 1.9 to 4.7 times speedup during the training phase and a 2.1 to 6.4 times increase in inference speed.
-Notably, the scalability concerns become even more evident as the size of the dataset expands; several graph-based models cannot complete the validation and testing processes within a reasonable time budget.
-The primary factors contributing to TNCN’s efficiency are the synergistic, time-efficient design of its two core components and the implementation of the Efficient CN Extractor that facilitates batch operations through parallel processing.
+Figure 3: Time Consumption of Memory and Graph-based Method on Wiki and Review Datasets.
+6.3 Ablation Study
 Reply to Q3: Extended CN brings improvements.
-Table 2: Test performance of TGN and TNCN with different ranges of common neighbors
+To elucidate the benefits of extended CN, we conducted an ablation study under official setting on the hop range of common neighbors.
+The results are shown in Table 2.
+Here we use notation "$k$-hop CN" to simply denote the CNs up to $(k,k)$-hop.
+The conventional NCN method considers only (1,1)-hop CN.
+However, this approach may not be universally applicable across all temporal networks.
+For instance, bipartite graphs lack such (1,1)-hop CN in their structure, necessitating the consideration of 2-hop CN.
+Additionally, memory-based methods may omit a notable aspect: they generally find it difficult to quantify the frequency of interactions between a given pair of nodes, which brings the need for 0-hop neighborhood.
+[…]
+Table 2: Test performance of TGN and TNCN with different ranges of common neighbors.
 Model Wiki Review Coin Comment
 TGN 0.528 0.387 0.737 0.622
-TNCN-1-hop-CN 0.621 0.374 0.737 0.641
-TNCN-0$\sim$1-hop-CN 0.720 0.298 0.739 0.692
-TNCN-0$\sim$2-hop-CN 0.717 0.317 0.764 0.662
-To elucidate the benefits of extended CN, we conducted an ablation study on the hop range of common neighbors.
-The results are shown in Table 2.
-The conventional NCN methodology considers only 1-hop CN.
-However, this approach may not be universally applicable across all temporal network topologies.
-For instance, bipartite graphs lack such 1-hop CN in their structure, necessitating the consideration of 2-hop CN.
-Additionally, memory-based methodologies may omit a notable aspect: they generally find it difficult to quantify the frequency of interactions between a given pair of nodes, which brings the need for 0-hop CN.
+TNCN-1-hop-CN 0.621 0.419 0.737 0.641
 […]
+[…]
+We have also conducted experiments for parameter analysis. Please refer to Appendix D.3 for details.
 7 Conclusion and Limitation
@@
 TNCN has achieved new state-of-the-art results on several real-world datasets while maintaining excellent scalability to handle large-scale temporal graphs.
-However, based on our observation of TNCN’s performance on the TGBL-Review dataset, there are some limitations in our model.
-Specifically, datasets with high surprise values, such as the Review dataset, tend to make it more challenging for TNCN to accurately predict the probability of future connections.
+However, based on our observation of TNCN’s performance on the Comment dataset, there are some limitations in our model.
+Specifically, non-bipartite datasets with high surprise values, such as the Comment dataset, tend to make it more challenging for TNCN to accurately predict the probability of future connections.
 This indicates that while TNCN performs well overall, it may struggle with datasets that exhibit high variability or unexpected patterns.
 Further research is needed to address these challenges and improve the model’s robustness in such scenarios.
+Acknowledgement
+This work is supported by the National Key R&\&D Program of China (2022ZD0160300) and National Natural Science Foundation of China (62276003).
 Appendix A Datasets
 Table 3 shows some detailed datasets statistics of TGB and 4 shows several temporal graph datasets commonly used by previous work.
-Through the two tables we can observe that TGB official datasets possess temporal graphs with larger scale to 10 million, 10 times surpassing the largest previous datasets such as LastFM, Flight, Contact.
-With the aim to examine our TNCN model’s efficiency, we choose the increasingly accepted datasets TGB.
-Table 3: TGB Dataset Statistics
+Through the two tables we can observe that TGB official datasets possess temporal graphs with larger scale to 10 million, 10 times surpassing the largest previous datasets such as LastFM.
+With the aim to examine our TNCN model’s efficiency, we choose the increasingly accepted datasets TGB in the main table.
+Table 3: TGB Dataset Statistics.
 Dataset Domain Nodes Edges Steps Surprise Edge Properties
@@
 tgbl-flight traffic 18143 67,169,570 1,385 0.024 W: $\times$, Di: ✓, A: ✓
-Table 4: Previous Dataset Statistics
-Datasets Domains Nodes Links NL Feat Bipartite Duration Unique Steps Time Granularity
+Here “Surprise index” [poursafaei2022towards] refers to the ratio of test edges that are not seen during training, which can be calculated as $\frac{|E_{test}/E_{train}|}{|E_{test}|}$.
+Low surprise index implies that memory-based methods such as Edgebank [poursafaei2022towards] may potentially achieve good performance, while high surprise may require more inductive capability.
+The surprise index varies across TGB datasets.
+Table 4: Previous Dataset Statistics.
+Datasets Domains Nodes Links N&L Feat Bipartite Duration Unique Steps Time Granularity
 Wikipedia Social 9,227 157,474 – & 172 $\checkmark$ 1 month 152,757 Unix timestamps
@@
 UCI Social 1,899 59,835 – & – $\times$ 196 days 58,911 Unix timestamps
-Flights Transport 13,169 1,927,145 – & 1 $\times$ 4 months 122 days
-UN Trade Economics 255 507,497 – & 1 $\times$ 32 years 32 years
-Contact Proximity 692 2,426,279 – & 1 $\times$ 1 month 8,064 5 minutes
-Appendix B TNCN Model Configuration
+Appendix B Related Work
+B.1 Memory-based Temporal Graph Representation Learning
+Temporal graph learning has garnered significant attention in recent years.
+A classic approach in this domain involves learning node memory using continuous events with non-decreasing timestamps.
+kumar2019predicting propose a coupled recurrent neural network model named JODIE that learns the embedding trajectories of users and items.
+Another contemporary work DyRep [trivedi2019dyrep] aims to efficiently produce low-dimensional node embeddings to capture the communication and association in dynamic graphs.
+rossi2020temporal introduces a memory-based temporal neural network known as TGN, which incorporates a memory module to store temporal node representations updated with messages generated from the given event stream.
+Apan [Wang_2021] advances the methodology by integrating asynchronous propagation techniques, markedly increasing the efficiency of handling large-scale graph queries.
+EDGE [chen2021efficient] emerges as a computational framework focusing on increasing the parallelizability by dividing some intermediate nodes in long streams each into two independent nodes while adding back their dependency by training loss.
+chen2023recurrent extend the update method for the node memory module, introducing an additional hidden state to record previous changes in neighbors.
+Complementing these efforts, additional contributions such as Edgebank [poursafaei2022towards] and DistTGL [zhou2023disttgl] have been directed towards formalizing and accelerating memory-based temporal graph learning methods.
+B.2 Graph-based Temporal Graph Representation Learning
+Subsequent works have incorporated the temporal neighborhood structure into temporal graph learning.
+CAWN [wang2021inductive] employs random anonymous walks to model the neighborhood structure.
+TCL [wang2021tcl] samples a temporal dependency interaction graph that contains a sequence of temporally cascaded chronological interactions.
+TGAT [xu2020inductive] considers the temporal neighborhood and feeds the features into a temporal graph attention layer utilizing a masked self-attention mechanism.
+NAT [luo2022neighborhood] constructs a multi-hop neighboring node dictionary to extract joint neighborhood features and uses RNN to recursively update the central node’s embedding.
+DyGFormer [yu2023towards], instead, leverages one-hop neighbor embeddings and the co-occurrence of neighbors to generate features, which are well-patched and subsequently fed into a Transformer [vaswani2017attention] decoder to obtain the final prediction.
+FreeDyG [tian2023freedyg] also utilizes historical interaction frequency akin to DyGFormer, afterwards transforming it with Fast Fourier Transform (FFT) and IFFT through the frequency domain.
+LPFormer [Shomer_2024] attempts to adaptively learn the pairwise encodings via graph attention module, utilizing relative position, ppr value and neighboring information to obtain the score.
+CNE-N [cheng2024co] uses a hash table to map an interaction event to its position.
+It calculates the co-neighbor encoding for each (neighbor - end node) pair within the local subgraph, recording the number of their common neighbors.
+These information are then concatenated to predict the probability of the future link.
+TPNet [lu2024improvingtemporallinkprediction] constructs temporal walk matrices via random propagation to simultaneously consider both temporal and structural information.
+Pairwise and auxiliary feature are then decoded to make the prediction.
+Another work RepeatMixer [zou2024repeatawareneighborsamplingdynamic] pays more attention to the repeat-aware neighbors.
+Such neighbor sequences are then leveraged and adaptively aggregated to learn the temporal patterns.
+DyG-Mamba [li2024dygmambacontinuousstatespace] introduces SSM to dynamic graph learning.
+It first extracts the first-hop interaction sequence between the given node pairs and encodes them.
+The time-span encoding between any two continuous timestamps are also computed, serving as control signals for the continuous SSM to obtain the final representation.
+CrossLink [crosslink10.1145/3696410.3714792] represents the graph evolution by a sequence of temporal events, where each event representation is obtained from a graph encoder.
+It finally utilizes a decoder-only transformer to model the sequence and predicts the next token, i.e.
+the link existence.
+EAGLE [li2025speedmeetsaccuracyefficient] aggregates the most recent neighbors of a central node and leverage temporal personalized PageRank value to capture the structural pattern, afterwards using adaptive weights to dynamically merge these features to get the prediction result.
+B.3 Link Prediction Methods
+Link prediction is a fundamental task in graph analysis, aiming to determine the likelihood of a connection between two nodes.
+Early investigations posited that nodes with greater similarity tend to be connected, which led to a series of heuristic algorithms such as Common Neighbors, Katz Index, and PageRank [newman2001clustering, katz1953new, page1999pagerank].
+With the advent of GNNs, numerous methods have attempted to utilize vanilla GNNs for enhancing link prediction, revealing sub-optimal performance due to the inability to capture important pair-wise patterns such as common neighbors [zhang2018link, zhang2021labeling, liang2022can].
+Subsequent research has focused on infusing various forms of inductive biases to retrieve intricate pair-wise relationships.
+For instance, SEAL [zhang2018link], Neo-GNN [yun2021neo], and NCN [wang2023neural] have integrated neighbor-overlapping information into their design.
+BUDDY [chamberlain2022graph] and NBFNet [zhu2021neural] have concentrated on extracting higher-order structural information.
+Additionally, mao2023revisiting, li2024evaluating have contributed to a more unified framework encompassing different heuristics.
+Appendix C TNCN Model Configuration
 Network Choice
 In our experiment, the changeable neural networks are chosen as follows:
-In Memory Module, we choose $Identity$ as msgfunc and GRU as $upd$.
+In Memory Module, we choose $Identity$ as msgfunc and GRU as updupd.
 In inference stage we process node memory with Graph Attention Embedding to get the temporal representation.
-As for Prediction Head, we finally choose $Linear$ as the $repr$ function.
+As for Prediction Head, we finally choose $MLP$ as the reprrepr function.
 Hyper-parameter
 Several detailed hyper-parameters for TNCN are shown in Table 5, which can help researchers to reproduce the experiment performance as reported in this paper.
-Additionally, we set $mem\_dim$, $emb\_dim$, $time\_dim$ 100, 100, 100 respectively.
-Table 5: Some Experiment Hyper-parameters
-Dataset num_neighbors num_epoch patience
-Wiki 10 20 5
-Review 10 5 2
-Coin 10 3 2
-Comment 10 3 1
-Appendix C Proofs
-In this section, we give proofs on theorems 5.4, 5.5 and 5.6.
-We commence with a restatement of Theorem 5.4 for clarity:
-Theorem C.1.
-(Ability of encoding $k$-hop event).
+Table 5: Some Experiment Hyper-parameters.
+Dataset num_neighbors num_epoch patience mem_dimmem\_dim emb_dimemb\_dim time_dimtime\_dim
+Wiki 15 20 5 184 184 100
+Review 15 10 3 184 184 100
+Coin 10 5 3 100 100 100
+Comment 10 3 2 100 100 100
+Appendix D Additional Experimental Results
+D.1 Transductive and Inductive Experiments on Previously Small and Medium Datasets
+In addition to the large-scale TGB dataset, we also conduct experiments on some traditional small and medium datasets previously used in dynamic graph link prediction.
+We follow TGN [rossi2020temporal] to evaluate different models under both transductive and inductive settings.
+The transductive setting deal with the future links between previously observed nodes in the training stage, and the inductive setting predicts link existence between unseen nodes.
+We compare TNCN with the aforementioned baselines with the addition of FreeDyG [tian2023freedyg], which is also a competitive method in temporal graph link prediction.
+The overall performance of TNCN and different models are in Table 6.
+From Table 6 we can find that TNCN achieves 6 new SOTA out of 7 datasets in transductive setting and 3 in inductive setting, furthur uncovering the strong performance of our model.
+Table 6: Average Precision (AP) under Transductive and Inductive settings on small and medium dataset.
+The best is in bold font, and the second is underlined.
+(“Trans” and “Ind” are the abbreviation for transductive and inductive respectively.)
+Setting Method Wikipedia Reddit Mooc Lastfm Enron Social Evo. UCI Avg. Rank
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+EdgeBank / / / / / / / /
+[…]
+[…]
+[…]
+[…]
+D.2 Comparison with Some Classic Heuristic Methods
+We exhibits the result between TGN with some classic heuristics and TNCN under official setting on tgbl-wiki dataset in the Table 7.
+Here heuristics consist of CN [doi:10.1126/science.286.5439.509], RA [zhou2009predicting], AA [adamic2003friends], PPR [page1999pagerank, jeh2003scaling] and ELPH [chamberlain2022graph].
+In these heuristic methods, the heuristic statistics are concatenated with TGN embedding to obtain final predictions.
+From the table we can see that these basic heuristics such as CN and RA do not bring performance improvement.
+However, some sophisticated heuristics like graph sketching in ELPH can enhance the backbone’s capability.
+Nevertheless, using these heuristics cannot outperform a more generalized model like our TNCN, which comprehensively takes neighborhood nodes’ representations into account.
+Table 7: Comparison between TGN with heuristics and TNCN on tgbl-wiki Dataset.
+Model Val MRR Test MRR Training Time (s) Inference Time (s)
+TGN 0.569 0.528 10.33 98.74
+TGN-CN 0.561 0.505 12.33 106.21
+TGN-RA 0.563 0.511 16.51 115.04
+TGN-AA 0.565 0.517 11.42 115.01
+TGN-PPR 0.521 0.427 207.01 327.22
+TGN-ELPH 0.715 0.681 240.92 1614.86
+TNCN 0.742 0.724 21.45 250.49
+D.3 Detailed Statistics of Parameter Analysis
+Here we provide some results on parameter analysis of our TNCN model.
+Figure 4 (a) illustrates the model performance under different numbers of neighbors.
+Incorporating more recent neighbors can retain richer information, while it may bring about some irrelevant noise.
+We can find that the test mrr increases as num_neighbors rises from a small value, but it declines after a certain threshold.
+In Figure 4 (b), we examine how the embedding and memory dimension affect the final performance.
+A similar trend is observed for this parameter.
+On the other hand, both the training and test time consistently increase with the parameter values becoming larger.
+This phenomenon is revealed in the Tables 8 to 11.
+In Table 8 and 9, we show the detailed MRR and time consumption of our TNCN with different numbers of neighbors over Wiki and Coin dataset.
+In Table 10 and 11 we show how the embedding and memory dimension influence the final performance.
+To strike a balance between the model capability and time consumption, we finally select the intermediate and appropriate values.
+(a) num_neighbors
+(b) emb_dim & mem_dim
+Figure 4: Parameter Analysis on Wiki Dataset.
+Table 8: Performance metrics for different numbers of neighbors on Wiki dataset.
+num_neighbors 10 12 15 18 20
+Val MRR 0.7379 0.7408 0.7412 0.7362 0.7418
+Test MRR 0.7164 0.7193 0.7240 0.7187 0.7183
+training time (s) 26.35 27.17 29.49 30.32 31.56
+test time (s) 378.88 396.75 407.43 413.83 420.22
+Table 9: Performance metrics for different numbers of neighbors on Coin dataset.
+num_neighbors 5 8 10 12 15
+Val MRR 0.7492 0.7378 0.7430 0.7450 0.7406
+Test MRR 0.7687 0.7619 0.7701 0.7662 0.7601
+training time (s) 5936.19 6129.33 6406.00 6911.00 7529.01
+test time (s) 56605.45 57117.99 57292.00 57745.00 57925.00
+Table 10: Performance metrics for different embedding and memory dimensions on Wiki dataset.
+emb_dim&mem_dim 100 150 184 256 512
+Val MRR 0.7402 0.7373 0.7412 0.7426 0.7404
+Test MRR 0.7178 0.7159 0.7240 0.7224 0.7235
+training time (s) 22.34 25.47 29.49 32.81 34.95
+test time (s) 378.28 399.77 407.43 420.45 459.86
+Table 11: Performance metrics for different embedding and memory dimensions on Coin dataset.
+emb_dim&mem_dim 30 50 100 150 184
+Val MRR 0.7387 0.7405 0.7430 0.7436 0.7518
+Test MRR 0.7591 0.7606 0.7701 0.7646 0.7699
+training time (s) 6228 6340 6406 6617 6721
+test time (s) 56694 57179 57292 58103 59411
+D.4 Comparison between TNCN and traditional NCN
+Here we provide a detailed comparison between our TNCN model and the traditional NCN method in Table 12.
+Table 12: The comparison of TNCN and NCN.
+temporal scenario backbone arbitrary CN hops batch-wise CN extraction
+NCN traditional GNN
+TNCN memory-based
+Appendix E Detailed Time and Memory Consumption Statistics
+E.1 Time Consumption
+Table 13 exhibits the detailed time consumption on TGB datasets with different models.
+We can observe that TNCN maintains similar time consumption to memory-based networks while achieving striking speedup compared with graph-based models.
+Table 13: Time Consumption of different methods on TGB Datasets. All these experiments are conducted with NVIDIA GeForce RTX 4090.
+Model(tr/val/test)(s) Wiki Review Coin Comment Flight
+TGN 12/41/43 988/1389/1411 2578/5219/5408 6599/9984/10311 2990/19834/20627
+DyGFormer 85/6268/6317 3891/25831/26911 12593/50893/51276 OOM OOM
+GraphMixer 27/2385/2425 983/14188/12764 5917/71699/72103 OOM OOM
+CNE-N 19/422/424 540/4749/4931 2308/18226/19317 OOM OOM
+DyRep 22/27/29 1514/877/916 6305/3821/3746 13701/6419/6587 9352/12188/13112
+TGAT 81/6359/6407 2564/33437/34814 25514/73372/73821 OOM OOM
+CAWN 173/15690/15881 5692/79566/80717 OOM OOM OOM
+TCL 31/680/682 948/9093/9962 4383/52973/54016 OOM OOM
+TNCN 33/565/566 971/2626/2701 5926/54765/55177 9828/59439/60163 8209/50533/51064
+Comparison between TNCN and NAT Here we also show some experimental results in Table 14 for the comparison between TNCN and NAT model.
+The hardware we use is NVIDIA GeForce RTX 2080 as NAT’s code isn’t compatible with higher version.
+Note that NAT model exposes a backward as its instability, accomplishing about only 1/3 experiments when we test it.
+Table 14: Comparison of Time Consumption between TNCN and NAT.
+Dataset Model Train (s) Val (s) Test (s)
+tgbl-wiki TNCN 21.45 250.49 251.52
+NAT 74.92 298.6 298.41
+tgbl-review TNCN 1649 4788 4695
+NAT 422 7516 7461
+tgbl-coin TNCN 4920 28716 28805
+NAT 1896 30398 30176
+E.2 Memory Consumption
+In Table 15 we have shown the memory cost of different methods on TGB datasets.
+On most of them (except Review) we can observe that TNCN takes similar GPU memory consumption to memory-based networks while occupying significant less memory than graph-based models.
+Table 15: Memory Consumption of different methods on TGB Datasets. All these experiments are conducted with NVIDIA GeForce RTX 4090. (“-” stands for out of memory)
+Inference Stage(MB) Wiki Review Coin Comment Flight
+TGN 1050 6416 11496 13838 1182
+DyGFormer 3174 5512 17428 - -
+GraphMixer 4914 5388 17360 - -
+CNE-N 1154 4628 16760 - -
+DyRep 998 6416 9300 10418 1126
+TGAT 5600 6588 18526 - -
+CAWN 7656 10898 - - -
+TCL 2204 5276 17292 - -
+TNCN 1420 9586 11990 13842 1188
+E.3 Overall Comparison between Performance and Time/Memory Consumption
+Here we provide a scatter plot to compare different methods between the final performance and the time/memory consumption.
+Figure 5 (a) shows the result on Wiki, and Figure 5 (b) on Coin.
+From these scatter plots, we can obtain a more intuitive perception that TNCN can better achieve a balance between the final performance and time/memory consumption.
+(a) Wiki
+(b) Coin
+Figure 5: Comparison between Performance and Time/Memory Consumption.
+Appendix F Pseudocode of TNCN Pipeline and CN Extraction
+Algorithm 1 shows the pseudocode about the pipeline of our TNCN model, and Algorithm 2 for the procedure of the CN extraction operation.
+Algorithm 1 Pipeline of TNCN
+1:for positive batch data $(\text{posu},\text{posv},t)$ do
+2:  $\text{neg\_batch}\leftarrow\text{negative sampling}$
+3:  for batch_data in {pos_data, neg_data} do
+4:   $\text{mem},\text{hist\_events}\leftarrow\text{memory\_module(batch\_data)}$;
+[…]
+6:   $\text{emb}\leftarrow\text{transform(mem, hist\_events)}$;
+[…]
+8:   $\text{CN\_mat}\leftarrow\text{CN\_extractor(hist\_events)}$;
+[…]
+10:   $\text{NCN\_emb}\leftarrow\text{AGG(emb, CN\_mat)}$;
+[…]
+12:   $p\leftarrow\text{Pred(emb\_u, emb\_v, NCN\_emb)}$;
+[…]
+14:  end for
+15:  $\text{mem}\leftarrow\text{memory\_update(pos\_data)}$;
+[…]
+17:end for
+Algorithm 2 Procedure of CN Extraction
+1:Input: Temporal adjacency matrix of node $u$ and $v$, denoted as $A\in\mathbb{R}^{N\times N}$ and $B\in\mathbb{R}^{N\times N}$.
+AA and $B$ are the sub-matrix from the whole adjacency matrix of the temporal graph, containing $u$, $v$ and their individual neighbors.
+NN is the total number of their temporal neighbors and themselves.
+2:Output: $k_{th}$ hop Common Neighbors of $u$ and $v$, $CN_{k}\in\mathbb{R}^{N}$.
+3:
+4:Stage 1: Generate up to $k$-hop neighbors.
+5:  $A\leftarrow A+I$, $B\leftarrow B+I$. $\triangleright$ add self-loop
+6:  Compute $A^{k-1}$ and $A^{k}$, $B^{k-1}$ and $B^{k}$. $\triangleright$ obtain up to k1k-1 / $k$ hop adjacency matrix
+7:  $N_{0\sim k-1}(u)=A^{k-1}[id(u)]$, $N_{0\sim k}(u)=A^{k}[id(u)]$. $\triangleright$ Similar to $v$. $N(u)\in\mathbb{R}^{N}$
+8:
+9:Stage 2: Get $0\sim k-1$ and $0\sim k$ hop common neighbors.
+10:  $_{0\sim k-1}=N_{0\sim k-1}(u)\odot N_{0\sim k-1}(v), \triangleright CNNC$$\in\mathbb{R}^{N}$
+11:  $CN_{0\sim k}=N_{0\sim k}(u)\odot N_{0\sim k}(v)$. $\triangleright$ perform sparse matrix hadamard product
+12:
+13:Stage 3: Get exact $k$-hop common neighbors.
+14:  $CN_{k}=CN_{0\sim k}-CN_{0\sim k-1}$. $\triangleright$ perform sparse matrix subtraction
+15:
+16:Finally we get the exact $k$-hop common neighbors between node $u$ and $v$.
+Appendix G Special Cases Analysis of Common Neighbor Extraction
+Here are some special cases while calculating $(1,2)$, $(2,1)$ and $(2,2)$ hop CNs.
+Under these situations, utilizing $A^{k}[id(u)]$ naively in step (2) will lead to walk-based neighbors, i.e.
+$\exists v,\exists i\neq j,w_{i}=w_{j},\ s.t.\ (u,w_{1},w_{2},\cdots,w_{k-1},v)\ exists.$ To obtain a clear version of arbitrary path-based $(i,j)$-hop CNs, we should avoid the repetition of neighbors.
+We take $(1,2)$ as an example to analyse the detailed method to eliminate repetition.
+Cases like $(2,1)$ and $(2,2)$ hop can be similarly solved.
+Assume that node $x$ is a $(1,2)$-hop CN of pair $(u,v)$, thus we know $\exists w,\ s.t.$ $(u,x)$ and $(v,w,x)$ exist.
+There are two variants that render $x$ to be a walk-based CN instead of a path-based one that we exactly require.
+(a) $x=v$.
+When $x=v$, the local graph has the topology shown in igure 6 a.
+This situation should satisfy two conditions: $w$ is a neighbor of $v$ and there are historical interactions between $u$ and $v$.
+Denote the frequency between $(u,v)$ before time $t$ as $q^{t}(u,v)=|\{(u,v,t^{\prime})|t^{\prime}<t\}\cup\{(v,u,t^{\prime})|t^{\prime}<t\}|$.
+So the naively computed $CN^{t}_{(1,2)}(u,v)[id(x)]$ value need to be subtracted by $\sum\limits_{w_{i}}q^{t}(w_{i},v)]*q^{t}(u,v)$, i.e.
+the total interaction frequency of $v$ before time $t$ multiplied by the frequency between $(u,v)$.
+(b) $w=u$.
+The structure is exhibited in Figure 6 b.
+Here $(u,v)$ has historical edges and $x$ is a $1$-hop neighbor of $u$.
+The additive substraction value is $\sum\limits_{x}q^{t}(x,u)^{2}]*q^{t}(u,v)$.
+(c) Both (1a) and (1b) are satisfied. The ground truth is as Figure 6 c. We just need to add back the overlap value that have been diminished once more.
+(a) $x=v$
+(b) $w=u$
+(c) $x=v$ & $w=u$
+Figure 6: Here shows the special cases related to $(1,2)$-hop CNs computation.
+Note that the graph is undirected, while the directed arrows implies the path direction used to determine the corresponding hop numbers.
+Note that the procedure above can only deal with CNs of no more than $(2,2)$-hop perfectly.
+For higher-order $(i,j)$-hop CN extraction, please refer to perepechko2009number for more details and complicated analysis.
+Appendix H Case Study
+H.1 Two Examples from TGB for Better Understanding TNCN’s Effectivity
+In Figure 7, we show two case studies from TGB to give a better understanding of the effectivity of our TNCN.
+(a) tgbl-wiki
+(b) tgbl-coin
+Figure 7: Two case studies from TGB.
+Figure (a) shows a case from tgbl-wiki, which is a bipartite graph.
+The yellow nodes 0 and 15 are a $(u,v)$ pair.
+If we use a node-wise method to predict the future link of $(0,15)$, we can find that node 15 has just 7 neighbors while node 0 has 12.
+So their properties may be different, thus having less chance to have an interaction.
+However TNCN can observe that the blue nodes are their $(1,2)$-hop CNs and purple nodes are the $(2,1)$-hop, and it will give a high probability over the existence of the future link.
+Figure (b) shows another case from tgbl-coin.
+Here we need to predict the link of $(1,8)$.
+Here TNCN can find that these two nodes have multiple variants of common neighbors.
+Node 3 is their $(1,2)$-hop CN, node 4 and 6 are the $(2,1)$-hop, and node 2 is both their $(1,1)$ and $(2,1)$ hop.
+The $(2,2)$-hop CNs are node 5 and 7, while node 9 being a special $(2,2)$-hop CN that owns a shared 1-hop node 2.
+With the abundant CN information, TNCN will be more likely to predict it as a positive future edge.
+H.2 Some Statistics about Common Neighbors on Comment Dataset
+Here we show some statistics about the structural information of common neighbors on Comment dataset, which may account for the lower performance of TNCN.
+We have collected four variants of statistics including (1) “s-nei/d-nei”: the number of recent neighbors (without repetition) of the positive src/dst nodes, (2) “cn”: the number of common neighbors of the positive node pairs, and (3) “prev”: their previous interaction times recently.
+All the values are the average from the test set.
+The statistics are shown in the following Table 16.
+Table 16: Statistics about CNs on Comment Dataset.
+(“s” stands for the “surprise” event, “ns” for not-surprise; “p” for the case where the ground-truth event ranks 1st in our prediction, “np” for not.)
+s-nei d-nei cn prev
+s-p 8.43 9.03 0.87 0.39
+s-np 8.89 8.78 0.07 0.02
+ns-p 7.83 8.05 2.64 0.69
+ns-np 8.76 8.88 0.17 0.07
+From the table we can find that, although src/dst nodes from both surprise and not-surprise edges have similar number of neighbors, their common neighbors and the previous contact frequecy are distinct far away.
+Nodes from surprising edges always share less CNs and interaction times, making CN features less effective in the prediction.
+From another aspect, the successfully predicted events typically have their nodes with more CNs and historical interactions, which indicates that our TNCN model may be misled by the node pair with more CNs in the prediction.
+Appendix I Proofs
+In this section, we give proofs on theorems 4.4, 4.7 and 4.6.
+Theorem I.1.
+(Theorem 4.4.
+Ability of encoding $k$-hop event).
 Given a $k$-hop event $\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$.
@@
 For the base case, $k=1$, the encoding has been shown to be feasible.
-Now, assume the proposition holds for a $k\textquoteright$ ’-hop event; that is, any $k\textquoteright$ ’-hop temporal sequence of monotonically increasing events can be encoded using a memory-based approach.
+Now, assume the proposition holds for a $k$’-hop event; that is, any $k$’-hop temporal sequence of monotonically increasing events can be encoded using a memory-based approach.
 This assumption implies that there exists an embedding function such that
-[…]
-[…]
+$Emb(u_{0},t_{u_{0},u_{1}})=\textit{Enc}({(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in{0,\ldots,k’-1}}),$ (7)
+for all event sequences with $k^{\prime}$ hops, where $k’\geq 1$.
+Given an arbitrary k+1k’+1-hop event, which can be partitioned into an initial event $(u_{0},u_{1},t_{u_{0},u_{1}})$ and a subsequent $k$’-hop sequence.
+The existence of an encoding function for the $k$’-hop sequence assures that
 $Emb(u_{0},t_{u_{0},u_{1}})$ $=f_{\textit{emb}}(Mem(u_{0},t_{u_{1},u_{2}}))$ (8)
-$=f_{\textit{emb}}(f_{\textit{mem}}(Mem(u_{0},t_{u_{1},u_{2}}),Mem(u_{1},t_{u_{1},u_{2}}),e_{u_{0},u_{1}}^{t_{u_{0},u_{1}}},t_{u_{0},u_{1}}-t_{u_{1},u_{2}}),$
+$=f_{\textit{emb}}(f_{\textit{mem}}(Mem(u_{0},t_{u_{1},u_{2}}),$
+$\quad Mem(u_{1},t_{u_{1},u_{2}}),e_{u_{0},u_{1}}^{t_{u_{0},u_{1}}},t_{u_{0},u_{1}}-t_{u_{1},u_{2}}),$
 $Mem(u_{1},t_{u_{1},u_{2}})$ $=f_{\textit{emb}}^{-1}(Emb(u_{1},t_{u_{1},u_{2}}))$
-$=f_{\textit{emb}}^{-1}(\textit{Enc}(\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})\mid i\in\{1,\ldots,k^{\prime}\},k^{\prime}\geq 1\})$
-[…]
+$=f_{\textit{emb}}^{-1}(\textit{Enc}(\{(u_{i},u_{i+1},t_{u_{i},u_{i+1}})$
+$\mid i\in\{1,\ldots,k^{\prime}\},k^{\prime}\geq 1\})$
+Subsequently, it is demonstrated that $Emb(u_{0},t_{u_{0},u_{1}})$ provides an encoding for both the initial event and the $k$’-hop sequence, thereby affirming its efficacy in encoding the entire k+1k’+1-hop event.
+This concludes the inductive step and substantiates the inductive argument.
 We consider a $k$-hop-subgraph-based approach for our analysis.
-It is evident that a $k$-hop subgraph encompasses any $k\textquoteright$ ’-hop events, where $k\textquoteright\leq k$.
+It is evident that a $k$-hop subgraph encompasses any $k$’-hop events, where $k’\leq k$.
 Furthermore, the aggregation methodology assimilates all nodes contained within the subgraph.
 Collectively, these observations substantiate the theorem in question.
-We commence with a restatement of Theorem 5.5 for clarity:
-Theorem C.2.
-(Learning method time complexity).
+Theorem I.2.
+(Theorem 4.7.
+Learning method time complexity).
 Denote the time complexity of a learning method as a function of the total number of events processed during training.
@@
 Consequently, the computational complexity is fundamentally proportional to $\sum_{u\in{\mathcal{N}}}d(u)^{2}$.
-Drawing on the results of de Caen [1998], the lower bound on the time complexity is established as $\Omega\left(\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}\right)$, whereas the upper bound is determined as ${\mathcal{O}}\left(\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}+|{\mathcal{E}}||{\mathcal{N}}|\right)$
-In the context of $k$-hop-subgraph-based algorithms wherein $k=2$,, we adopt similar strategy where each event$(u,v,t)$ will only be utilized once another event within the subgraph of $u$ or $v$ is firstly considered.
+Drawing on the results of DECAEN1998245, the lower bound on the time complexity is established as $\Omega\left(\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}\right)$, whereas the upper bound is determined as ${\mathcal{O}}\left(\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}+|{\mathcal{E}}||{\mathcal{N}}|\right)$
+In the context of $k$-hop-subgraph-based algorithms wherein $k=2$, we adopt similar strategy where each event(u,v,t)(u,v,t) will only be utilized once another event within the subgraph of $u$ or $v$ is firstly considered.
 The total number of events can be formulated as $\sum_{u\in{\mathcal{N}}}d(u)\sum_{v\in{\mathcal{N}}_{u}}\sum_{w\in{\mathcal{N}}_{v}}d(w)$.
@@
 Following Cauchy inequality and conclusions of $\sum_{u\in{\mathcal{N}}}d(u)^{2}$, we got the the upper-bound time complexity is ${\mathcal{O}}\left((\frac{|{\mathcal{E}}|^{2}}{|{\mathcal{N}}|}+|{\mathcal{E}}||{\mathcal{N}}|)^{\frac{3}{2}}\right)$
-We commence with a restatement of Theorem 5.6 for clarity:
-Theorem C.3.
+Theorem I.3.
+(Theorem 4.6. Expressivity of TNCN)
+  1. 1.
     TNCN is strictly more expressive than CN, RA, and AA.
-We first give definitions of these structural features under temporal settings. Given two nodes $u$ and $v$, the structural features before time $t$ are defined as follows:
+  2. 2.
+    TNCN is strictly more expressive than Jodie with the same dimension of time encoding, DyRep with the same aggregation function, TGAT with the same attention layers and neighbors, and TGN under identical condition for all module choices.
+Proof.
+(1) We first give definitions of these structural features under temporal settings. Given two nodes $u$ and $v$, the structural features before time $t$ are defined as follows:
 $CN(u,v,t)$ $=\sum_{w\in N_{1}^{t}(u)\cap N_{1}^{t}(v)}1,$ (9)
@@
 Given a node $u$, the degree of node $u$ is the number of events $e$ with an endpoint at node $u$.
-Without loss of generality (WLOG), we consider node $u$ as the source node, and the events are $\{(u,v_{i},t_{i})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$.
+Without loss of generality (W.L.O.G.), we consider node $u$ as the source node, and the events are $\{(u,v_{i},t_{i})\mid i\in\{0,\ldots,k-1\},k\geq 1\}$.
 Each time a new event is given, the embedding of node $u$ is updated by
@@
 $emb_{u}^{t}=f(d(u)).$ (12)
-Thus, the neural common neighbor $TNCN_{1}(u,v)=\mathop{\oplus}_{w\in N_{1}^{t}(u)\cap N_{1}^{t}(v)}emb_{w}^{t}$ can express Equation 9.
+Thus, the neural common neighbor
+$TNCN_{1}(u,v)=\mathop{\oplus}_{w\in N_{1}^{t}(u)\cap N_{1}^{t}(v)}emb_{w}^{t}$ can express Equation 9.
 Extending to situations where the common neighbor node has some features we want to learn, the traditional CN, RA, and AA cannot accommodate this.
 However, our TNCN can express these features, demonstrating that TNCN is strictly more expressive than CN, RA, and AA.
+(2) We first prove that TNCN is strictly more expressive than TGN.
+As TNCN possesses the same memory-based framework as TGN to derive the $emb_{u}^{t}$, TNCN at least has the same expressivity as TGN with identical module choices.
+Additionally, from part (1) of this theorem we know that TNCN can capture certain heuristics such as CN where TGN fails.
+Hence TNCN is strictly more expressive than TGN.
+Since Jodie, DyRep and TGAT are specific instances of TGN with minor changes, we next shift our object to prove that TGN has no less expressivity than the three methods.
+For TGAT, it retains only the Graph Attention mechanism for embedding functions, removing all memory and message modules present in TGN.
+Consequently, TGAT lacks the capacity to incorporate temporal memory updates and message passing, making it less expressive than TGN.
+Jodie differs from TGN in the memory updater and embedding function.
+Specifically, it utilizes an RNN rather than a GRU, although GRU is a general variant of RNN.
+Additionally, Jodie employs the time embedding as $emb_{u}^{t}=(1+\Delta t\cdot w)\otimes mem_{u}^{t^{-}}$, where $w$ are learnable parameters.
+However, TGN’s Graph Attention uses not only the memory $mem_{u}^{t^{-}}$ but also the historical edge features $e_{u,v}^{t^{\prime}}$, and learns the coefficient via multi-head attention, thus achieving a higher expressivity.
+DyRep also modifies the memory updater to an RNN.
+Meanwhile, it changes the embedding function to Identity and message function to Graph Attention.
+Although it just swaps these two modules, the expressivity is reduced in the temporal setting.
+When the model attempts to make its prediction, it first generates the embeddings with the embedding function.
+So naturally the identity function processes less information in the $emb_{u}^{t}$ than graph attention.
+After obtaining the prediction result, the new messages will be calculated by the message function and fed into memory updater, where the $emb_{u}^{t}$ can not get access.
+Hence TGN is more expressive than DyRep, especially when the recent information can make large benefits.
+In conclusion, TNCN is strictly more expressive than Jodie, DyRep, TGAT and TGN under the same condition.